### v1

# Online Coreset Selection for Learning Dynamic Systems

Abstract
With the increasing availability of streaming data in dynamic systems, a critical challenge in data-driven modeling for control is how to efficiently select informative data to characterize system dynamics. In this work, we design an online coreset selection method under the framework of set-membership identification for systems subject to process disturbances, with the objective of improving data efficiency while ensuring convergence guarantees. Specifically, we first propose a stacked polyhedral representation that over-approximates the feasible set of system parameters. Leveraging a generalized Grünbaum’s inequality, we design a geometric selection criterion for constructing the coreset. To reduce computational complexity, an online double-description-based constraint reduction method is introduced to simplify the polyhedral representation. Finally, we analyze the convergence of the feasible set with respect to the coreset and derive upper bounds on the selection probability and the expected number of data in the coreset. The effectiveness of the proposed method is demonstrated through comprehensive simulation studies.
{IEEEkeywords}
Online coreset selection; Set-membership identification; Convergence analysis; Polyhedral computing;
1 Introduction
[…]
Recent years have seen rapid advancements in data-driven control (DDC) [1, 2, 3]. Among various methodologies for DDC, a well-established procedure involves first characterizing the set of system models consistent with observed data, often termed the feasible set, and subsequently designing a robust controller to satisfy control objectives for all systems within this set [3]. The former task belongs to set-membership identification (SMI), which aims to learn a nominal set of systems from data [4]. In online learning processes, a major challenge arises from the fact that dynamic systems continuously generate a huge volume of data. Directly applying all raw data for DDC will lead to an increasing computational burden and unlimited storage growth [5]. Consequently, developing efficient strategies to enhance data efficiency while preserving essential information is crucial for practical deployment of set-based data-driven methodologies in real-time control tasks.
Coreset selection has been developed as a prominent technique to efficiently processing large dataset, particularly in machine learning and data analytics [6, 7, 8]. The objective is to identify a small yet representative subset of the original dataset so that models or parameters learned from this coreset closely approximate those derived from the full dataset. This approach has demonstrated notable success in accelerating various tasks on static, large-scale datasets, such as supporting vector machine [9], K-clustering [10], logistic regression [11, 12], Gaussian mixture models [13, 14] and deep learning [15, 16]. Nevertheless, extending coreset selection method to dynamic systems introduce significant challenges primarily due to the inherent online nature, the violation of the independent and identically distribution (i.i.d.) assumption, and the task-specific goals of system identification and control. Despite these difficulties, recent efforts in DDC have explored intelligent data selection strategies to improve data efficiency for dynamic systems. For example,[17] proposed an event-triggered mechanism for Gaussian processes, where data are selected based on model uncertainty. In[18], data selection is triggered when prediction error or distance exceeds a threshold under a Lazily Adapted Constant Kinky Inference model. In [19], new data are incorporated when the system state deviates significantly from a reference trajectory, prompting model updates. However, for fully-parameterized linear systems with bounded disturbance noises, the design of coreset selection strategies that ensure provable performance guarantees remains an open challenge.
To enhance data efficiency in dynamic system learning, SMI provides a natural and powerful framework. Instead of deriving point estimates (e.g., least squares methods), SMI characterizes the entire set of parameters consistent with observed data and prior disturbance bounds [20, 4, 21], offering a principled way to assess the informativeness of new data. Several methods for enhancing data efficiency in online learning process have been explored under this framework. One line of research is called bounded complexity set update, which constrains the feasible set’s complexity – e.g., via polytopes with fixed hyperplane directions [22], parallelotopes [23], or compact over-approximations [24]. Another line of works focus on informativeness-driven updates, where the selecting data that are benefits for reducing uncertainty. For instances, the feasible set is updated only when incoming data sufficiently reduce uncertainty characterized  [4, 25]. While these approaches aim to improve data efficiency, a crucial aspect remains underexplored: do such selection schemes guarantee convergence of the feasible set? This question is especially critical when representing feasible set from limited or irregularly sampled data streams.
With the above discussions, the convergence of feasible set in online learning processes emerges as a critical concern. This form of convergence exhibits two key challenges that distinguish it from standard parameter convergence analysis. First, it corresponds to worst-case convergence of the points in a set. Second, the convergence is concerned under subsets of entire data. The analysis on the worst-case convergence under the entire dataset has been explored in many previous works [26, 27, 28, 29, 30, 31]. In [26], a point-wise asymptotic convergence of the feasible set for discrete-time scalar system was derived under tight noise and persistently exciting regressors. [27] investigated the convergence of size of feasible set. For linear systems, [28] derived an asymptotic convergence property under persistent excitation of regressors and tight noise. In [31], a non-asymptotic convergence of size of feasible set under tight noise and block-martingale small-ball (BMSB) condition was proved. The convergence properties under irregular or event-triggered sampling have also attracted increasing attention [32, 33, 34]. In [32], an event-triggered binary identification algorithm was proposed with guaranteed convergence. In [33], convergence of parameters is discussed considering multi trajectories using least squares methods, and the convergence rate is derived. Similarly, [34] approached learning from irregularly-sampled time series from a missing data perspective, proposing generative models within an encoder-decoder framework to capture the underlying data distributions. However, the convergence properties of feasible set under a selected coreset have not been analyzed before and needs further exploration.
In this paper, we study the problem of online coreset selection design for learning unknown systems within a set-membership identification framework. We focus on linear time-invariant system with bounded additive disturbances, and aim to design a criterion for selecting data in online learning processes. The goal is to construct a coreset that yields a feasible set with significantly fewer data points, while still ensuring convergence guarantees for the learning algorithm. To achieve this goal, we first propose an over-approximated polyhedral representation of feasible set for general bounded disturbance descriptions. Inspired by the work of generalized Grunbaum’s inequality [35], we design a selection criterion based on the geometrical relation between new data points and existing uncertainty representation. However, several challenges need to be addressed to enable our design. First, how can the feasible set be effectively represented using only the coreset data? Moreover, how to update the polyhedral representation as new data points are selected. In addition, does the dataset selection criterion based on the geometrical relation guarantees the convergence? How to derive the asymptotic and non-asymptotic expression of size of feasible set. The main contributions of this work are summarized as follows:
  1. 1.
    An explicit stacked polyhedral representation for the feasible system set is proposed. Specifically, we derive a full-parameter characterization of all feasible system under bounded disturbance, where each row of the uncertain system matrix lies within a polyhedral set defined by available data. The proposed representation captures the entire feasible system in a computationally tractable and geometrically convenient form and is applicable to general bounded disturbance models (Proposition 1).
  2. 2.
    An online coreset selection criterion is designed based on the geometric structure of the uncertainty set. Leveraging the derived polyhedral representation, a threshold-triggered selection mechanism is designed to determine whether a newly arrived data point should be included in the coreset. This mechanism relies on the geometric relationship between new hyperplane constraint and the current uncertainty set. To improve computational efficiency, an online double-description method is adopted to simplify the update of polyhedral representation (Algorithm 2).
  3. 3.
    Theoretical guarantees on the convergence of the feasible set are established (Theorems 4.134.25). We show that the volume of the feasible polytope decreases to zeros with respect to (w.r.t.) selected coreset. Both asymptotic and non-asymptotic convergence results are derived based on generalized Grunbaum’s inequality and persistent excitation on process data. Furthermore, we provide probabilistic bounds on selection and derive an upper bound for expected cardinality of coreset (Theorem 4.28).
The rest of the paper is organized as follows. Section 2 introduce preliminaries and problem formulation. The polyhedral representation of feasible set, the design of the online coreset selection criterion and the double-description based constraints reduction are presented in Section 3. The main theoretic results, including the convergence of feasible set constructed from coreset, as well as upper bounds on selection probability and the coreset cardinality are provided in Section 4. Several numerical examples and implementation results are represented and discussed in Section 5, followed by some concluding remarks in Section 6.
Notation: Let $\mathbb{R},\mathbb{Z},\mathbb{R}_{\geq 0},\mathbb{Z}_{>0}$ denote the set of real numbers, integers, nonnegative real numbers and positive integers respectively. For integers $i_{1}<i_{2}$, let $[i_{1}:i_{2}]\triangleq\{i_{1},i_{1}+1,\dots,i_{2}\}$ and let $[N]\triangleq[1:N]$. Let $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times m}$ denote the $n$-dimensional Euclidean space and the space of $n\times m$ real matrices, respectively. Let $\emptyset$ denote the empty set. For $x,y\in\mathbb{R}^{n}$, let $\langle x,y\rangle$ denote the inner product, and $x(i)$ its $i$-th coordinate. Let $\mathbf{1}_{n}\triangleq[1,\dots,1]^{T}\in\mathbb{R}^{n}$ be the all-ones vector. For $x,y\in\mathbb{R}^{n}$, write $x<y$ if $x(i)<y(i)$ for all $i\in[n]$, and $x\leq y$ analogously. For $M\in\mathbb{R}^{n\times m}$, let $M(i,j)$, $M(i,:)$, and $M(:,j)$ denote the $(i,j)$-th entry, $i$-th row, and $j$-th column, respectively. $M_{1}\succeq M_{2}$ denotes that $M_{1}-M_{2}$ is positive semi-definite. Denote the Kronecker product by $M_{1}\otimes M_{2}$ and the Cartesian product of $m$ copies of $\mathbb{R}^{n}$ by $(\mathbb{R}^{n})^{m}$. For $x\in\mathbb{R}$, let $|x|$ denote the absolute value. For $x\in\mathbb{R}^{n}$, let $\|x\|_{2}=\sqrt{\langle x,x\rangle}$ denote the 2-norm of vector. For a finite set $\mathcal{A}$, $|\mathcal{A}|$ denotes its cardinality. The convex and conic hulls of the columns of a matrix $K$ are denoted by $\text{conv}(K)\triangleq\{K\lambda\mid\lambda\geq 0,\,\sum_{i}\lambda_{i}=1\}$ and $\text{cone}(K)\triangleq\{K\eta\mid\eta\geq 0\}$, respectively. For sets $X,X^{\prime}\subseteq\mathbb{R}^{n}$, the Minkowski sum is $X+X^{\prime}\triangleq\{x+x^{\prime}\mid x\in X,\,x^{\prime}\in X^{\prime}\}$.
2 Preliminaries and Problem Formulation
Consider a discrete linear time-invariant system
$x_{k+1}=Ax_{k}+Bu_{k}+w_{k},$ (1)
where $x_{k}\in\mathbb{R}^{n_{x}}$ is the state of the system at time $k$, $u_{k}\in\mathbb{R}^{n_{u}}$ is the control input, $w_{k}\in\mathbb{R}^{n_{x}}$ is the bounded disturbance, $A\in\mathbb{R}^{n_{x}\times n_{x}}$ and $B\in\mathbb{R}^{n_{x}\times n_{u}}$ are unknown system matrices.
Let $z_{k}\triangleq\begin{bmatrix}x_{k}^{T}&u_{k}^{T}\end{bmatrix}^{T}\in\mathbb{R}^{n_{z}}$ be a state-input vector at time $k$ where $n_{z}=n_{x}+n_{u}$. For any $K>0$ and any set $S=\{k_{1},...,k_{|S|}\}\subseteq[K+1]$, we denote $Z_{S}\triangleq[z_{k_{1}},\dots,z_{k_{|S|}}]$ and $X_{S}\triangleq[x_{k_{1}},\dots,x_{k_{|S|}}]$ as grouped state-input matrix and state matrix indexed by $S$, respectively.
2.1 Preliminaries
2.1.1 Set-membership Identification
In this section, we review SMI for linear system with bounded noise. Consider the system described by equation (1). Suppose that the additive disturbance $w_{k}$ satisfies the following assumption:
Assumption 1 (Bounded stochastic disturbance)
The disturbance sequence $\{w_{k}\}_{k\geq 0}\subset\mathbb{R}^{n_{x}}$ consists of independent and identically distributed (i.i.d.) random vectors, each following a probability measure $\mathbb{P}$ supported on a compact, convex set $\mathcal{W}\subset\mathbb{R}^{n_{x}}$. The set $\mathcal{W}$ is assumed to be symmetric w.r.t. the origin.
Remark 1
Several existing works propose specific structural assumptions on the disturbance set $\mathcal{W}$. For instance, $\mathcal{W}$ is modeled as a bounded polytope in [36], an ellipsoid in [37], and a zonotope in [38]. In contrast, this paper considers a more general setting by assuming only that $\mathcal{W}\subset\mathbb{R}^{n_{x}}$ is a known compact and convex set. This allows for a broader class of uncertainty models while retaining tractability in the subsequent analysis.
Let $\Theta_{*}=\begin{bmatrix}A&B\end{bmatrix}$ denote the true system parameters. For any subset $\mathcal{S}\subseteq\mathbb{Z}_{>0}$, we define
$\Theta_{\mathcal{S}}\triangleq\bigcap_{k\in\mathcal{S}}\{{\Theta}:x_{k}-{\Theta}z_{k-1}\in\mathcal{W}\}.$ (2)
Then, by construction, the true system parameters $\Theta_{*}$ satisfies:
$\Theta_{*}\in\Theta_{\mathcal{S}}.$ (3)
The intersection $\Theta_{\mathcal{S}}$ provides a set-valued characterization of the uncertainty in the system parameters, which can be viewed as an uncertainty qualification for $\Theta_{*}$. The right-hand side of (3) contains all feasible parameters consistent with the input-output data.
2.1.2 Polyhedra and Polytopes
We introduce some basic concepts and properties related to polyhedral representation in this section. Let $H\in\mathbb{R}^{n\times m}$ be a matrix and $c\in\mathbb{R}^{n}$ be a vector. A polyhedron in $\mathcal{H}$-representation (also called an $\mathcal{H}$-polyhedron) is defined by
$P_{\mathcal{H}}(H,c)\triangleq\{x\in\mathbb{R}^{m}\mid Hx\leq c\},$ (4)
which is the solution set of a system of linear inequalities.
Correspondingly, let $V_{1}\in\mathbb{R}^{m\times n_{v}}$ and $V_{2}\in\mathbb{R}^{m\times n_{r}}$ be real matrices. A polyhedron in $\mathcal{V}$-representation (also called a $\mathcal{V}$-polyhedron) is defined as the Minkowski sum of the convex hull of columns of $V_{1}$ and conic hull of columns of $V_{2}$, i.e.,
$P_{\mathcal{V}}(V_{1},V_{2})\triangleq\text{conv}(V_{1})+\text{cone}(V_{2}).$ (5)
The Minkowski-Weyl theorem [39] states that a subset $P$ in $\mathbb{R}^{m}$ has an $\mathcal{H}$-representation if and only if there exist matrices $V_{1}$ and $V_{2}$ such that $P=P_{\mathcal{V}}(V_{1},V_{2})$. Thus, every polyhedron has two representations of type $P_{\mathcal{H}}$ and $P_{\mathcal{V}}$.
A polytope is a bounded polyhedron. From the definition of $\mathcal{V}$-polyhedron, the conic term is absent in this case, i.e., $\text{cone}(V_{2})$ is an empty set. In this case, we write
$P_{\mathcal{V}}(V_{1})\triangleq\text{conv}(V_{1})$ (6)
to denote the resulting bounded polyhedron.
The centroid of a polytope $P$ is defined as
[…]
where $\mu(\cdot)$ denotes the Lebesgue measure of an $m$-dimensional convex body.
The supporting function $h_{P}:\mathbb{R}^{m}\to\mathbb{R}$ for a convex set $P\subset\mathbb{R}^{m}$ is
$h_{P}(\xi)\triangleq\max_{x\in P}\{\langle x,\xi\rangle\}.$ (8)
We denote the centered supporting function $\widetilde{h}_{P}:\mathbb{R}^{m}\to\mathbb{R}$ w.r.t. the centroid $g(P)$ as :
$\widetilde{h}_{P}(\xi)\triangleq\max_{x\in P}\{\langle x-g(P),\xi\rangle\}.$ (9)
Then, by [40], we have the following inequalities:
$\frac{1}{m}\widetilde{h}_{P}(\xi)\leq\widetilde{h}_{P}(-\xi)\leq m\widetilde{h}_{P}(\xi).$ (10)
2.2 Problem Formulation
Due to uncertainty of $w$, we cannot obtain $\Theta_{*}$ precisely from dataset. Alternatively, we leverage set-membership identification to characterize the feasible set of $\Theta_{*}$ and design selection mechanism in an online fashion to enhance data efficiency. Over time period $[K]$, if all process data are kept as coreset (the trivial case), the SMI of $\Theta_{*}$ by (2) is represented by $\Theta_{[K]}$. Denote $\mathcal{D}_{k}=\{z_{k-1},x_{k}\}$ as the dataset to be determined to select at time step $k$ which determines a feasible set as shown in (2). Let $\mathcal{S}_{K}\subset[K]$ be a subset of all time steps until $K$. Then, the coreset constructed for SMI is represented by
$\mathcal{D}_{\mathcal{S}_{K}}=\cup_{k\in\mathcal{S}_{K}}\mathcal{D}_{k}.$ (11)
The corresponding SMI for $\Theta_{*}$ is denoted by $\Theta_{\mathcal{S}_{K}}$.
Our objective is to design an online coreset selection criterion to update $\mathcal{S}_{K}$. The selected coreset yields a reduced dataset whose size is significantly smaller than total number of samples $K$, while ensuring some learning performance guarantees. However, designing such a selection criterion is nontrivial due to several intertwined challenges:
First, the basic SME formulation (2) does not explicitly characterize the estimation uncertainty, i.e., it does not reveal how the volume or size of the feasible set evolves as new data (e.g., $x_{k+1}$) arrive. An analytical representation of the feasible set is required for general $\mathcal{W}$. This motivates the first problem:
Problem 1: How can we construct a tractable and analytical representation of the feasible set that enables quantification and comparison of estimation uncertainty?
Given such a representation, the next challenge is to develop a principled efficient data selection criterion that determines whether newly arriving data should be retained.
Problem 2: How to design a selection criterion that retains informative data points for reducing uncertainty of feasible set?
Finally, and most critically, the designed selection criterion should preserve the theoretical guarantees of the SMI framework. In particular, it is required that the feasible set constructed in an online fashion using coreset converges to true parameters as more data become available. This leads to the third problem:
Problem 3: Under the proposed coreset selection criterion, how to establish convergence properties of the feasible set? What are the statistical properties of the selection probability and the cumulative number of selected data points in online learning processes?
3 Online Coreset Selection Design
In this section we address Problems 1 and 2. First, we derive a stacked polyhedral over-approximation of feasible parameters set, which provides a tractable and analytical representation of the estimation uncertainty. Subsequently, we propose a coreset selection mechanism that identifies and retains informative data points. Finally, based on the double description method for polyhedra, we develop a dynamic constraints reduction algorithm to further optimize the representation by eliminating redundant constraints.
3.1 Polyhedral Over-approximation of Feasible Set
The main idea behind deriving the data-based representation of full parameterized feasible set is to decompose uncertain matrix $\Theta\subset\mathbb{R}^{n_{x}\times n_{z}}$ into $n_{x}$ row vectors in $\mathbb{R}^{n_{z}}$, each constrained within a data-dependent polyhedron. To illustrate the idea, we denote row-stacking map $\Phi:(\mathbb{R}^{n})^{m}\to\mathbb{R}^{m\times n}$ as
$\Phi(v_{1},...,v_{m})=\begin{bmatrix}v_{1}^{T}&\dots&v_{m}^{T}\end{bmatrix}^{T},$ (12)
where each $v_{i}\subset\mathbb{R}^{n}$ is treated as a column vector set, and $v_{i}^{T}$ denotes its transpose(a row vector set). Then, matrix set $\Theta\subset\mathbb{R}^{n_{x}\times n_{z}}$ can be represented as $\Theta=\Phi(\theta^{1},\dots,\theta^{n_{x}})$ where $\theta^{i}\subset\mathbb{R}^{n_{z}}$ for all $i\in[n_{x}]$. The true system parameters $\Theta_{*}$ can be written as $\Theta_{*}=\Phi(\theta_{*}^{1},\theta_{*}^{2},\dots,\theta_{*}^{n_{x}})$ where $\theta_{*}^{i}\in\mathbb{R}^{n_{z}}$ are constant vectors. Denote initial uncertainty matrix set $\Theta_{0}$ as
$\Theta_{0}\triangleq\Phi(\theta_{0}^{1},\dots,\theta_{0}^{n_{x}}).$ (13)
We consider the following assumptions on $\Theta_{0}$:
Assumption 2 (Initial Admissible Set)
$\forall i\in[n_{x}]$, $\theta_{*}^{i}$ lies in a known, bounded, and convex polytope $\theta_{0}^{i}=P_{\mathcal{H}}(H_{0}^{i},c_{0}^{i})$, i.e., $\theta_{*}^{i}\in\theta_{0}^{i}=\{\theta^{i}:H_{0}^{i}\theta^{i}\leq c_{0}^{i}\}$.
Note that assumptions on prior bounded uncertainty are common in set-membership identification [36]. Now we are ready to derive a data-driven representation of uncertainty set. Let $\bar{w}\in\mathbb{R}^{n_{x}}$ be a vector whose entries are the component-wise maxima of $\mathcal{W}$. That is
$\forall w\in\mathcal{W},w\leq\bar{w}.$ (14)
Due to symmetry assumption of $\mathcal{W}$, we have $-\bar{w}\leq w\leq\bar{w}$. For $i\in[n_{x}]$, take
$P(\mathcal{D}_{[K]}^{i})\triangleq\{\theta\in\mathbb{R}^{n_{z}}:H_{K}^{i}\theta\leq C_{K}^{i}\}$ (15)
to be a polyhedral set, where
$H_{K}^{i}\triangleq\begin{bmatrix}H_{0}^{i}\\ Z_{[0:K-1]}^{T}\\ -Z_{[0:K-1]}^{T}\end{bmatrix},C_{K}^{i}\triangleq\begin{bmatrix}c_{0}^{i}\\ (X_{[K]}(i,:))^{T}+\overline{W}(i)\\ -(X_{[K]}(i,:))^{T}+\overline{W}(i)\end{bmatrix},$ (16)
$\overline{W}(i)\triangleq\mathbf{1}_{K}\otimes\bar{w}(i).$ (17)
The following proposition provides a data-driven representation of the uncertainty set.
Proposition 1
Consider system (1) with full dataset $\mathcal{D}_{[K]}$. Suppose Assumptions 1 and 2 hold. Then, the stacked polyhedral set
$\widehat{\Theta}_{K}^{\text{full}}\triangleq\Phi(P(\mathcal{D}_{[K]}^{1}),P(\mathcal{D}_{[K]}^{2}),...,P(\mathcal{D}_{[K]}^{n_{x}}))$ (18)
satisfies
$\widehat{\Theta}_{K}^{\text{full}}\supseteq(\Theta_{0}\cap\Theta_{[K]}),$ (19)
where $\Theta_{0}$ and $\Theta_{K}$ are defined in (13) and (3), respectively.
Proof 3.1.
We complete the proof by showing that $\widehat{\Theta}_{K}^{\text{full}}$ is an over-approximation of $\Theta_{0}\cap\Theta_{[K]}$. For all $w\in\mathcal{W}$, every coordinate of $w$ is bounded by $\bar{w}$. Then, we have:
$\begin{bmatrix}I\\ -I\end{bmatrix}w\leq\begin{bmatrix}\bar{w}\\ \bar{w}\end{bmatrix}.$ (20)
Based on SME, for each $\Theta=\Phi(\theta^{1},\dots,\theta^{n_{x}})\in\Theta_{0}\cap\Theta_{[K]}$, we have $x_{k}-\Theta z_{k-1}\in\mathcal{W}$. Thus,
$\begin{bmatrix}I\\ -I\end{bmatrix}(x_{k}-\Theta z_{k-1})\leq\begin{bmatrix}\bar{w}\\ \bar{w}\end{bmatrix}$ (21)
holds for all $k\in[K]$. For all $i\in[n_{x}]$, stacking the $i$-th and the $(i+n_{x})$-th rows horizontally and transposing the resulting matrix, we infer from (21) that
$\begin{bmatrix}z_{k-1}^{T}\\ -z_{k-1}^{T}\end{bmatrix}\theta^{i}\leq\begin{bmatrix}x_{k}(i)+\bar{w}(i)\\ -x_{k}(i)+\bar{w}(i)\end{bmatrix}.$ (22)
Then, by incorporating the initial feasible constraints of $\theta^{i}$ in Assumption 2, we have $\theta^{i}\in P(\mathcal{D}_{K}^{i})$ holds for all $i\in[n_{x}]$. Consequently, $\Theta\in\widehat{\Theta}_{K}^{\text{full}}$, and thus $\widehat{\Theta}_{K}^{\text{full}}\supseteq(\Theta_{0}\cap\Theta_{[K]})$.
Remark 3.2.
Proposition 1 provides an explicit data-based representation $\widehat{\Theta}_{K}^{\text{full}}$ of the true parameters $\Theta_{*}$ by SME using the full dataset over time period $[K]$. Compared with existing polyhedral representation of feasible set that directly operate on $\mathrm{vec}(\Theta_{*})\in\mathbb{R}^{n_{x}n_{z}}$(i.e. [41]), the proposed representation has a key advantage: each polyhedral component of $\widehat{\Theta}_{K}^{\text{full}}$ lies in a significantly lower-dimensional space $\mathbb{R}^{n_{z}}$. This dimensionality reduction enables the use of standard polyhedral operations—such as conversions between $\mathcal{V}$- and $\mathcal{H}$-representations— for systems of moderate size, which could reduce computational burden commonly encountered in high-dimensional polyhedral set representations. By this representation, one can achieve a favorable trade-off: although the number of polyhedra increases, their individual low dimension renders many otherwise intractable operations computationally feasible.
3.2 Criterion for Coreset Selection
In this subsection we focus on the problem of how to design coreset selection criterion for data selection and thus improve data efficiency.
We suppose that the learning process has continued up to time step $K-1$. The set $\Phi(P_{K-1}^{1},P_{K-1}^{2},\dots,P_{K-1}^{n_{x}})$ is a data-based representation containing true parameters $\Theta_{*}$ where $P_{K-1}^{i}$ for $i\in[n_{x}]$ are polyhedral sets constructed from selected process data. At time instant $K$, for each $i\in[n_{x}]$, the sampling of $x_{K}(i)$ generates two half-space constraints which can be written as
$H_{+}^{i}$ :=$:=\{\theta:\quad(z_{K-1})^{T}\theta\leq x_{K}(i)+\bar{w}(i)\},$ (23)
$H_{-}^{i}$ :=$:=\{\theta:(-z_{K-1})^{T}\theta\leq-x_{K}(i)+\bar{w}(i)\}.$ (24)
These half-space constraints can be equivalently expressed via centered supporting function. For the half-space constraints $H_{+}^{i}$ in (23), it can be reformulated as:
$-z_{K-1}^{T}(\theta-g(P_{K-1}^{i}))\geq-x_{K}(i)-\bar{w}(i)+z_{K-1}^{T}g(P_{K-1}^{i}).$ (25)
The right-hand side of (25) can be rewritten as $\alpha_{+}^{i}\widetilde{h}_{P_{K-1}^{i}}(z_{K-1})$. Here, $\widetilde{h}_{P_{K-1}^{i}}(z_{K-1})$ denotes the centered supporting function defined in (9), and the scaler $\alpha_{+}^{i}$ is given by
$\alpha_{+}^{i}=\frac{-x_{K}(i)-\bar{w}(i)+z_{K-1}^{T}g(P_{K-1}^{i})}{\widetilde{h}_{P_{K-1}^{i}}(z_{K-1})}.$ (26)
We can apply the same procedure to $H_{-}^{i}$ and obtain the analogous expression:
$H_{-}^{i}=\{\theta:z_{K-1}^{T}(\theta-g(P_{K-1}^{i}))\geq\alpha_{-}^{i}\widetilde{h}_{P_{K-1}^{i}}(-z_{K-1})\}$ (27)
with the corresponding scaling parameter:
$\alpha_{-}^{i}=\frac{x_{K}(i)-\bar{w}(i)-z_{K-1}^{T}g(P_{K-1}^{i})}{\widetilde{h}_{P_{K-1}^{i}}(-z_{K-1})}.$ (28)
Then, we have the following proposition in a compact form.
Proposition 3.3.
Let $K\in\mathbb{Z}_{>0}$ be fixed. For each $i\in[n_{x}]$, the constraints $H_{+}^{i}$ and $H_{-}^{i}$ in (23) and (24) admit the equivalent representation in terms of $\alpha_{+}^{i}$ and $\alpha_{-}^{i}$ in the following form:
$H_{+}^{i}$ =$=\{\theta:\langle-z_{K-1},\theta-g(P_{K-1}^{i})\rangle\geq\alpha_{+}^{i}\widetilde{h}_{P_{K-1}^{i}}(z_{K-1})\},$
$H_{-}^{i}$ =$=\{\theta:\langle z_{K-1},\theta-g(P_{K-1}^{i})\rangle\geq\alpha_{-}^{i}\widetilde{h}_{P_{K-1}^{i}}(-z_{K-1})\},$
where $\alpha_{+}^{i}$ and $\alpha_{-}^{i}$ are given in (26) and (28) respectively.
Remark 3.4.
The scaling parameters $\alpha_{+}^{i}$ and $\alpha_{-}^{i}$ quantify the normalized penetration depth of the hyperplanes into the polytope. Consider half-space $H_{+}^{i}$ and the polytope $P_{K-1}^{i}$ as an example: When $\alpha_{+}^{i}=-1$, (23) can be reformulated as $z_{K-1}^{T}(\theta-g(P_{K-1}^{i}))\leq\widetilde{h}_{P_{K-1}^{i}}(z_{K-1})$. This inequality holds for all $\theta\in P_{K-1}^{i}$. Equality is achieved when $\theta=\arg\max_{x\in P_{K-1}^{i}}\{\langle x-g(P_{K-1}^{i}),z_{K-1}\rangle\}$, indicating that $H_{+}$ is tangent to $P_{K-1}^{i}$ at its supporting hyperplane. When $\alpha_{+}^{i}=0$, (23) can be reformulated as $z_{K-1}^{T}(\theta-g(P_{K-1}^{i}))\leq 0$, indicating that the hyperplane $\{\theta:z_{K-1}^{T}\theta=x_{K}(i)+\bar{w}(i)\}$ passes exactly through the centroid $g(P_{K-1}^{i})$. For $\alpha_{+}^{i}\in(-1,0)$, it’s guaranteed that $H_{+}^{i}\cap P_{K-1}^{i}\neq\emptyset$ and $H_{+}^{i}\cap P_{K-1}^{i}\subset P_{K-1}^{i}$, which shows that adding ${z_{K-1},x_{K}^{i}}$ at time instant $K$ can decrease size of feasible set.
Based on the above discussion, we propose the following selection criterion for constructing the coreset.
Let $\alpha_{0}\in[-1,0)$ be a predefined threshold. For each dimension $i\in[n_{x}]$ and time instant $K$, define the sub-trigger $\gamma_{K,i}\in\{0,1\}$ as
$\gamma_{K,i}=1\iff\max(\alpha_{+}^{i},\alpha_{-}^{i})\geq\alpha_{0},$ (29)
where $\alpha_{+}^{i},\alpha_{-}^{i}$ are determined in (26) and (28) respectively. Let $\vee$ denotes the logical OR, the binary trigger $\gamma_{K}\in\{0,1\}$ is defined as :
$\gamma_{K}=1\iff\vee_{i=1}^{n_{x}}\gamma_{K,i}=1.$ (30)
The coreset is constructed recursively according to
$\mathcal{D}_{K}=\begin{cases}\mathcal{D}_{K-1}\cup\{z_{K-1},x_{K}\}&\text{if }\gamma_{K}=1,\\ \mathcal{D}_{K-1}&\text{if }\gamma_{K}=0.\end{cases}$ (31)
The trigger $\gamma_{K}=1$ implies that there exists $i\in[n_{x}]$ such that a newly constructed half-space constraints determined by $z_{K-1}$ and $x_{K}^{i}$ has a scaling parameter $\alpha^{i}$, with respect to $P_{K-1}^{i}$, which is larger than $\alpha_{0}$. Depending on the choice of $\alpha_{0}$, two types of selection rules can be defined: When $\alpha_{0}=-1$, the selection trigger in (30) is termed as immediate trigger, since any intersection between the hyperplane and the uncertainty polytope will instantaneously activate the triggering condition. Also, for $\alpha_{0}\in(-1,0)$, we term the selection trigger in (30) as threshold trigger, since only those hyperplanes with scaling parameter larger than $\alpha_{0}$ activate triggering. The feasible set of $\Theta_{*}$ can be represented by $\Theta_{0}\cap\Theta_{\mathcal{S}_{K}}$.
3.3 Dynamic Double Description Constraint Reduction
The mechanism defined in (2931) determines the coreset to be retained during the online learning process. However, for all $i$ with $\gamma_{K,i}=1$, directly incorporating both $H_{+}^{i}$ and $H_{-}^{i}$ into the polyhedral set construction for uncertainty representation may lead to redundant descriptions. To reduce redundancy, we propose a double description-based constraint reduction method in this section. A complete update scheme for the uncertainty set is also presented.
The online double description-based constraint reduction process can be formalized as follows: Given a polyhedral set $P$ with non-redundant $\mathcal{H}$-representation $P_{\mathcal{H}}(H_{0},C_{0})$ and $P_{\mathcal{V}}(V_{0})$ and a new half-space constraints: $\{h^{T}x\leq c\}$, the goal is to obtain the updated non-redundant $\mathcal{H}$-representation $P_{\mathcal{H}}(H,C)$ and $\mathcal{V}$-representation $P_{\mathcal{V}}(V)$ of the resulting new polyhedral set. The pseudocode of the proposed constraint reduction method is presented in Algorithm 1.
Algorithm 1 Double Description-based Constraint Reduction for Dynamic polyhedra(DDCR-DYN)
0:  Initial polytope $P_{\mathcal{H}}(H_{0},C_{0})$ and $P_{\mathcal{V}}(V_{0})\subset\mathbb{R}^{n}$, a hyperplane-vertex logical matrix $G_{0}$, a new hyperplane constraint $\{h^{T}x\leq c\}$, a calculation precision $\varepsilon$.
0:  $H,C,V,G$
1:  Homogenization: $\widetilde{V_{0}}\leftarrow[1_{|V_{0}|}^{T};V_{0}]$, $\widetilde{H_{0}}\leftarrow\begin{bmatrix}-C_{0}&H_{0}\end{bmatrix}$
2:  $\begin{bmatrix}\alpha_{1}&\dots&\alpha_{|V_{0}|}\end{bmatrix}=\begin{bmatrix}-c&h^{T}\end{bmatrix}\widetilde{V_{0}}$
3:  $J_{+}=\{i:\alpha_{i}\geq\varepsilon\}$, $J_{-}=\{i:\alpha_{i}<-\varepsilon\}$
4:  $J_{o}=\{i:-\varepsilon\leq\alpha_{i}\leq\varepsilon\}$
5:  if $J_{+}=\phi$ then return $H_{0},C_{0},V_{0},G_{0}$
6:  $\widetilde{V}=\widetilde{V_{0}}$, $H=[H_{0};h^{T}]$, $C=[C_{0};c]$.
7:  $G=[G_{0}(H_{0},V_{0});\mathbf{0}]$.
8:  for $i\in J_{-},j\in J_{+},\text{Adjacent}(v_{i},v_{j})$ do
9:     $\tilde{v}=\alpha_{j}\tilde{v_{i}}-\alpha_{i}\tilde{v_{j}}$
10:     $\bar{H}(v)=\{\bar{h}:\bar{h}^{T}\tilde{v}\geq-\varepsilon,\bar{h}^{T}\in\text{row}(\widetilde{H_{0}})\}$
11:     Normalize $\tilde{v}$ as $\tilde{v}/\tilde{v}(1)$
12:     Update $\widetilde{V}=[\widetilde{V}\,\tilde{v}]$ and $G=[G\quad(\widetilde{H_{0}}\tilde{v}\geq-\varepsilon)]$
13:  end for
14:  Remove $\tilde{v}_{i}\in\widetilde{V}$ if $i\in J_{+}$ and update $V=\widetilde{V}(2:end,:)$.
15:  Remove $h\!\in\!\text{row}(H)$ if $G(h,:)\!==\!0$ and update $H,C,G$
16:  return $H,C,V,G$
Now we explain steps in Algorithm 1 in detail. The inputs to the Algorithm 1 are an initial polytope with both $\mathcal{H}$-representation $P_{\mathcal{H}}(H_{0},C_{0})$ and $\mathcal{V}$-representation $P_{\mathcal{V}}(V_{0})$, a hyperplane-vertex logic matrix $G_{0}$, a new hyperplane constraint $\{h^{T}x\leq c\}$ and a minimal calculation precision $\varepsilon$. For a matrix $V$, let $\text{row}(V)$ denote set $\{v:v\text{ is a row vector in }V\}$ and $\text{col}(V)$ denote set $\{v:v\text{ is a column vector in }V\}$. The hyperplane-vertex logic matrix $G_{0}$ is defined by
$G_{0}(i,j)=\begin{cases}1&v_{i}\in\text{col}(V_{0}),h_{j}^{T}\in\text{row}(H_{0}),h_{j}^{T}v_{i}=c_{j},\\ 0&\text{else}.\end{cases}$ (32)
Therefore, $G_{0}(i,j)=1$ can be geometrically interpreted by the vertex $v_{i}$ lying on hyperplane $\{x:h_{j}^{T}x=c_{j}\}$. The output of DDCR-Dyn is a polytope constituting original polyhedral constraints and new hyperplane constraint while maintaining minimal representation.
Lines 1-11 are similar to the classical double description method: We first homogenize the vertex $v$ as $\tilde{v}=[1;v]$ and the half-space constraints becomes $[-c,h][1;x]\leq 0$. Then we derive set $J_{+},J_{-}$ and $J_{o}$ by those $\alpha$ calculated in line 2. For all adjacent $v_{i}$ and $v_{j}$ where $i\in{J_{-}}$ and $j\in J_{+}$, we calculate new vertices in line 9 and add them to set $\widetilde{V}$ after normalization. The key procedure in Algorithm 1 is to update logical matrix $G$. For better illustration of its update, we consider an example of constraints reduction in 2D in Fig. 1.
Figure 1: Dynamic update of logical matrix $G$ during constraint reduction. Left: a 2D polyhedron with a newly added hyperplane $h_{\text{new}}$, causing redundancy in the original $h_{3}$. Right: Evolution of matrix G: Gray regions denote vertices in $J_{+}$(violating $h_{\text{new}}$ constraints), while the dashed box highlights the absence of active vertices for $h_{3}$ after removal of $J_{+}$.
The left part of Fig. 1 represents an existing 2D polygon $P(\mathcal{D}_{S_{k-1}}^{i})$ (dark gray region) and new hyperplane constraints $h_{\text{new}}$ (dashed blue line) determined by $h$ and $c$. During the addition of $h_{\text{new}}$, vertices in $J_{+}$ are marked in light gray. Those generated vertices satisfying condition $\bar{H}(v)=1$ ($v_{6}$ and $v_{7}$ in Fig. 1) are kept and added to active vertices set. To remove redundancy (Line 14 and 15 in Algorithm 1), first we remove those vertices $v_{i}$ with $i\in J_{+}$ and update $G$. Then, all redundant hyperplanes $h$ are removed by looking up to $G$ and finding all zeros rows. In our example, the hyperplane $h_{3}$ becomes a redundant hyperplane, with $G(h_{3},v)=0$ for all $v$, which is shown in dashed box in Fig. 1.
Let $\mathcal{D}_{k}^{i}\triangleq\{z_{k-1},x_{k}(i)\}$. The complete online learning procedure, including coreset selection and uncertainty set update, is shown in Algorithm 2. At each time instant $k$, upon receiving a new data point $x_{k}$, the data selection criterion (30) is tested. If $\gamma_{k}=1$, the data $\mathcal{D}_{k}$ is retained as part of the coreset. For each coordinate direction $i$ such that $\gamma_{k,i}=1$, the polyhedral set $P_{k}^{i}$ is updated via the DDCR-DYN subroutine: $P_{k}^{i}=f_{\text{DDCR-dyn}}(P_{k-1}^{i},\mathcal{D}_{k}^{i})$. The updated feasible set $\widehat{\Theta}_{k}$ is then reconstructed from the polyhedral set $P_{k-1}^{i}$ using the mapping $\Phi(\cdot)$. If no update is triggered (i.e., $\gamma_{k}=0$), the feasible set remains unchanged: $\widehat{\Theta}_{k}=\widehat{\Theta})_{k-1}$.
The function $f_{\text{DDCR-dyn}}(P_{k-1}^{i},\mathcal{D}_{k}^{i})$ in Algorithm 2 implements a two-stage sequential process: it first takes the initial $\mathcal{H}$ and $\mathcal{V}$-representation of $P_{k-1}^{i}$ and half-space constraints $H_{+}^{i}$ as the input of DDCR-DYN. Then, at the second stage, it uses the output of the first stage as the new input polytope and processes the second half-space constraints $H_{-}^{i}$.
Algorithm 2 Online Learning
0:  Double description of initial feasible set $\Theta_{0}$: $H_{0},C_{0}$ and $V_{0}$, a maximum calculation precision $\varepsilon$, a triggering threshold $\alpha_{0}$.
0:  Coreset $\mathcal{D}_{k}$ and feasible set $\widehat{\Theta}_{k}$ at each time $k$.
1:  for $k=1,2,\dots$ do
2:     for $i=1,\dots,n_{x}$ do
3:        Calculate $\alpha_{+}^{i}$ and $\alpha_{-}^{i}$ by (26) and (28).
4:        Determine $\gamma_{k,i}$ by (29) with triggering threshold $\alpha_{0}$.
5:     end for
6:     Calculate $\gamma_{k}$ by (30) and get $\mathcal{D}_{k}$ by (31)
7:     if $\gamma_{k}==1$ then
8:        for $i=1,\dots,n_{x}$ do
9:           If $\gamma_{k,i}=1$, $P_{k}^{i}=f_{\text{DDCR-dyn}}(P_{k-1}^{i},\mathcal{D}_{k}^{i})$.
10:        end for
11:        $\widehat{\Theta}_{k}=\Phi(P_{k-1}^{1},\dots,P_{k-1}^{n_{x}})$
12:     else
13:        $\widehat{\Theta}_{k}=\widehat{\Theta}_{k-1}$.
14:     end if
15:  end for
4 Convergence Analysis
In this section, we analyze the convergence of the feasible set constructed using the coreset selected by method in Section 3. First, we discuss the convergence of measure of feasible set under the threshold trigger. Then, the convergence properties under the immediate trigger are analyzed. Finally, the selection probability bound and the related cumulative number of selected data points for the proposed online coreset selection mechanism are discussed.
4.1 Convergence under Threshold Trigger $(\alpha_{0}\in(-1,0))$
In this section, we discuss the convergence of feasible set under general threshold trigger condition. We begin by introducing the following two assumptions on additive disturbance $w_{k}$ and state-input vector $z_{k}$, respectively.
The disturbance distribution $\mathbb{P}$ introduced in Assumption 1 satisfies the following condition:
Assumption 3 (Tight Disturbance Bounds)
There exist two non-decreasing functions $q_{w},Q_{w}:\mathbb{R}_{\geq 0}\to[0,1]$, with $q_{w}(0)=Q_{w}(0)=0$ and $q_{w}(\epsilon)>0$ for all $\epsilon>0$, such that for all $\tilde{w}\in\partial\mathcal{W}$ and $\epsilon>0$, the following two-sided inequality holds:
$q_{w}(\epsilon)\leq\mathbb{P}\left(w\in\mathcal{W}:\|w-\tilde{w}\|_{2}<\epsilon\right)\leq Q_{w}(\epsilon).$ (33)
The state-input vector $z_{k}$ satisfies the well-known persistent excitation condition:
Assumption 4 (Persistent excitation)
There exists $\beta>0$, $N_{u}>0$ and $b_{z}>0$ such that ${||z_{k}||}_{2}\leq b_{z}$ and for all $k_{0}>0$, the following holds:
$\frac{1}{N_{u}}\sum_{k=k_{0}+1}^{k_{0}+N_{u}}z_{k}z_{k}^{T}\succeq\beta^{2}I.$ (34)
Remark 4.5.
Assumptions 3-4 are consistent with standard assumptions commonly used in the literature (e.g., [26, 31, 27]). The key difference is the explicit inclusion of an upper bound function $Q_{w}(\cdot)$, which characterizes how concentrated $\mathbb{P}$ can be near the boundary. This addition does not strengthen the assumption, as the existence of such a bound is guaranteed by the fact that $\mathbb{P}$ is a probability measure.
These assumptions serve as the foundation for our convergence analysis. We then introduce some lemma which are required to establish the convergence of feasible set under coreset. If the disturbance distribution $\mathbb{P}$ satisfies the two-sided inequality in (33), we say that $\mathbb{P}$ has the two-sided boundary-visiting property with respect to the functions $(q_{w},Q_{w})$. Accordingly, we refer to $q_{w}$ and $Q_{w}$ as the boundary functions of the additive disturbance $w$.
Let $\mathcal{W}_{i,\epsilon}\triangleq[-\bar{w}(i),-\bar{w}(i)+\epsilon)\cup(\bar{w}(i)-\epsilon,\bar{w}(i)]$, $i\in[n_{x}]$. The two-sided boundary-visiting property can be extended to coordinate-wise cases shown in the following lemma.
Lemma 4.6.
There exist $q:\mathbb{R}_{\geq 0}\to[0,1]$ and $Q:\mathbb{R}_{\geq 0}\to[0,1]$ such that $q(0)=Q(0)=0$, $q$ and $Q$ are non-decreasing and $q(\epsilon)>0$ for all $\epsilon>0$. The marginal measure $\mathbb{P}_{i}$ on $\pi_{i}(\mathcal{W})=[-\bar{w}(i),\bar{w}(i)]$ satisfies the two-sided boundary property w.r.t. $\mathit{q}$ and $\mathit{Q}$. Specifically, for any $\epsilon>0$,
$q(\epsilon)\leq\mathbb{P}_{i}\left(w\in\mathcal{W}_{i,\epsilon}\right)\leq Q(\epsilon).$ (35)
Proof 4.7.
The proof is provided in the Appendix 7.
Remark 4.8.
Lemma 4.6 indicates that each coordinates of $w$ has two-sided boundary visiting property under Assumption 3. This can be explained as there exists a uniform probability bound for each coordinate of $w$ from its boxed boundary.
Under Assumption 4, we have the following result.
Lemma 4.9.
Let Assumption 4 be satisfied. For all $\vartheta\in\mathbb{R}^{n_{z}}$ with $\left\lVert\vartheta\right\rVert_{2}=\delta>0$, and for any $k_{0}>0$, it holds that
$\max_{k\in[k_{0}+1:\,k_{0}+N_{u}]}|\vartheta^{T}z_{k}|\geq\beta\delta.$ (36)
Proof 4.10.
From Assumption 4, pre- and post-multiplying both sides of (34) by $\vartheta^{T}$ and $\vartheta$, and using $\|\vartheta\|_{2}=\delta$, we obtain
$\frac{1}{N_{u}}\sum_{k=k_{0}+1}^{k_{0}+N_{u}}|\vartheta^{T}z_{k}|^{2}\geq\beta^{2}\delta^{2}.$ (37)
By the pigeonhole principle, there exist at least one $k$ such that $|\vartheta^{T}z_{k}|\geq\alpha_{0}\delta$, which implies (36).
Next, we introduce a generalization of Grünbaum’s Inequality proposed in [35].
Lemma 4.11.
Let $P\subset\mathbb{R}^{n_{z}}$ be a convex body with centroid at origin. Let $\alpha\in(-1,0)$ and $\xi\in S^{n_{z}-1}$. Consider the half-space
$H_{\alpha}^{+}=\{x\in\mathbb{R}^{n_{z}}:\langle z,\xi\rangle\geq\alpha h_{P}(-\xi)\}.$ (38)
Then
$\frac{\mu({P\cap H_{\alpha}^{+}})}{\mu{(P)}}\leq C(\alpha,n_{z}),$ (39)
where
$C(\alpha,n_{z})=1-\left(\frac{n_{z}(\alpha+1)}{n_{z}+1}\right)^{n_{z}}.$ (40)
Here, the generalization allows the hyperplane to be shifted by a factor $\alpha\in(-1,0)$ along a given direction, thus accommodating more flexible partitioning.
Remark 4.12.
In the context of the proposed coreset selection strategy, this lemma quantifies the reduction in the measure (volume) of the feasible polyhedral set after each data selection. The parameter $\alpha_{+}^{i}$ and $\alpha_{-}^{i}$ in Proposition 4.19 correspond to the parameter $\alpha$ in (38). Therefore, the threshold trigger condition (29) can be equivalently understood as requiring that each selection is triggered only when the feasible set undergoes a guaranteed relative volume reduction, characterized by the factor $C(\alpha_{0},n_{z})$. Proposition 3.3 and Lemma 4.11 together offer a geometric explanation of the proposed online coreset selection strategy.
Given the row-stacking map $\Phi$ defined in (12), it is natural to define the measure of $\Phi(\cdot)$ via a product measure. Specifically, we consider product measure $\mu(\Phi(\cdot))$ satisfying
$\mu(\Phi(P_{1},\dots,P_{n_{x}}))=\prod_{i=1}^{n_{x}}\mu(P_{i}),$ (41)
where each $P_{i}$ is a polyhedral set of $\mathbb{R}^{n_{z}}$, and $\mu(P)(P\subset\mathbb{R}^{n_{x}})$ denotes the standard Lebesgue measure on $\mathbb{R}^{n_{z}}$, as previously used in the definition of the centroid in (7). We note that a slight abuse of notation is involved here: the symbol $\mu$ is used both for the Lebesgue measure on $\mathbb{R}^{n_{z}}$ (when applied to $P_{i}$) and for the product measure on $\mathbb{R}^{n_{x}n_{z}}$ (when applied to $\Phi(P_{1},\dots,P_{n_{x}})$), whose domain and dimensionality are different.
Let $n^{i}(K)$ denote the cumulative number of selected data points caused by the $i$-th component up to time $K$, that is,
$n^{i}(K)=\sum_{k=1}^{K}1\{\gamma_{k,i}=1\},$ (42)
where $1\{\cdot\}$ denotes the indicator function.
Denote by $K^{i}(n)$ the time step at which the $i$-th component triggers for the $n$-th time, with the convention $K^{i}(0)=0$. Further, for any $K\geq 0$, define $\tau^{i}(K)$ as the first time after $K$ when the $i$-th component triggers selection, i.e.,
$\tau^{i}(K)\triangleq\min\{k>K\mid\gamma_{k,i}=1\}.$ (43)
For a bounded polyhedral set $P$ and an inner point $\theta\in P$, we define the radius $\mathfrak{R}(P,\theta)$ as
$\mathfrak{R}(P,\theta)\triangleq\sup_{\theta^{\prime}\in P}\|\theta-\theta^{\prime}\|_{2}.$ (44)
Now we are in a position to state the theorem for convergence for measure of feasible set under threshold trigger selection condition.
Theorem 4.13.
Consider system (1) with threshold trigger $(\alpha_{0}\in(-1,0))$. Let Assumptions 1-4 be satisfied, and $q(\epsilon)$ be a lower bound function in (35). The proposed estimator $\widehat{\Theta}_{K}$ has the following convergence properties.
  • (i)
    The measure $\mu(\widehat{\Theta}_{K})$ is bounded as:
    $\mu(\widehat{\Theta}_{K})\leq\mu(\Theta_{0})\prod_{i=1}^{n_{x}}C(\alpha_{0},n_{z})^{n^{i}(K)},$ (45)
    where $C(\alpha_{0},n_{z})$, defined in (40), is independent of $K$.
  • (ii)
    For all $i\in[n_{x}]$, with probability no less than $1-p_{\varepsilon}$, the following holds:
    $\frac{\tau^{i}(K)-K}{N_{u}}\leq{\ln(p_{\varepsilon})}/{\ln(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right))},$ (46)
    where $p_{\varepsilon}>0$, $\delta=\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})$ and $\beta$ are defined in (34).
  • (iii)
    The following asymptotic convergence holds almost surely:
    $\lim_{K\to\infty}\mu(\widehat{\Theta}_{K})=0.$ (47)
Proof 4.14.
For each triggering event associated with the $i$-th polyhedral set, by Lemma 4.11, the measure of this feasible polytope is reduced by at least a factor of $C(\alpha_{0},n_{z})$. Since $n^{i}(K)$ denotes the number of triggering event for the $i$-th sub-trigger up to time $K$, it follows that
$\mu(P_{K}^{i})\leq\mu(\theta_{0}^{i})C(\alpha_{0},n_{z})^{n^{i}(K)}.$ (48)
Substituting the above bound in (41) yields
$\mu(\widehat{\Theta}_{K})\leq\prod_{i=1}^{n_{x}}\mu(\theta_{0}^{i})\,C(\alpha_{0},n_{z})^{n^{i}(K)}$ (49)
$=\mu(\Theta_{0})\prod_{i=1}^{n_{x}}C(\alpha_{0},n_{z})^{n^{i}(K)}.$ (50)
This establishes the desired inequality in statement (i).
Fix $i\in[n_{x}]$ and $K>0$. For $K_{2}\geq K_{1}>K$, we define event $\mathcal{E}_{[K_{1}:K_{2}]}$ as
$\mathcal{E}_{[K_{1}:K_{2}]}=\{\forall k\in[K_{1}:K_{2}],\gamma_{k,i}=0\}.$ (51)
To prove statement (ii), we need to derive an upper bound for $\mathbb{P}(\mathcal{E}_{[K+1:K+\Delta K]})$.
Consider $k\in[K+1:K+\Delta K]$. By (29), $\gamma_{k,i}=1$ if and only if $\alpha_{+}^{i}\geq\alpha_{0}$ or $\alpha_{-}^{i}\geq\alpha_{0}$. By (1), we have $x_{k}(i)=z_{k-1}^{T}\theta_{*}^{i}+w_{k-1}(i)$. Then, $\alpha_{+}^{i}\geq\alpha_{0}$ can be rewritten as
$w_{k-1}(i)\leq-\bar{w}(i)+\psi_{k-1}^{+},$ (52)
where
$\psi_{k-1}^{+}=\left((-\alpha_{0})\widetilde{h}_{P_{k-1}^{i}}(z_{k-1})+z_{k-1}^{T}(g(P_{k-1}^{i})-\theta_{*}^{i})\right).$ (53)
Similarly, the condition $\alpha_{-}^{i}\geq\alpha_{0}$ holds if and only if
$w_{k-1}(i)\geq\bar{w}(i)-\psi_{k-1}^{-},$ (54)
where
$\psi_{k-1}^{-}=\left((-\alpha_{0})\widetilde{h}_{P_{k-1}^{i}}(-z_{k-1})-z_{k-1}^{T}(g(P_{k-1}^{i})-\theta_{*}^{i})\right).$ (55)
Here, $\psi_{k-1}^{+}$ and $\psi_{k-1}^{-}$ are determined by the available information at time $k-1$ (i.e., $z_{k-1}$ and $P_{k-1}^{i}$), and is independent of the realization of $w_{k-1}$. Also, we denote $\psi_{k}\triangleq\max\{\psi_{k}^{+},\psi_{k}^{-}\}$.
Now we assume that the event $\mathcal{E}_{[K+1:K+\Delta K]}$ holds, i.e., $\gamma_{k,i}=0$ for all $k\in[K+1:K+\Delta K]$. Then the feasible set $P_{K}^{i}$ remains unchanged during this time interval. Let $\theta^{i}_{K}=\arg\sup_{\theta\in P_{K}^{i}}\|\theta_{*}^{i}-\theta\|_{2}$ and define the vector $\vartheta\triangleq\theta^{i}_{K}-\theta_{*}^{i}$, we have $\left\lVert\vartheta\right\rVert_{2}=\delta$.
Under Assumption 4, with $\alpha_{0}\in(-1,0)$, we show that $\forall k_{0}\in[K:K+\Delta K-N_{u}]$, the following expression holds:
$\max_{k\in[k_{0}+1:k_{0}+N_{u}]}\psi_{k}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta.$ (56)
By Lemma 4.9, denote $k^{\prime}={\arg\max}_{k\in[k_{0}+1:\,k_{0}+N_{u}]}|\vartheta^{T}z_{k}|$. Then, $|\vartheta^{T}z_{k^{\prime}}|\geq\beta\delta$. We first consider the case $z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})\geq 0$, and distinguish two sub-cases based on the sign of $\vartheta^{T}z_{k}$. (The case for $z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})<0$ will be discussed later in a similar process)
  • If $\vartheta^{T}z_{k}\geq\beta\delta$, invoking the definition of $\widetilde{h}_{P_{K}^{i}}(z_{k^{\prime}})$ in (9) and observing that $\theta_{K}^{i}\in P_{K}^{i}$ during the interval under consideration, we obtain
    $\psi_{k^{\prime}}^{+}$ $=(-\alpha_{0})\max_{\theta\in P_{K}^{i}}{\langle\theta-g(P_{K}^{i}),z_{k^{\prime}}\rangle}+z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})$
    $=(-\alpha_{0})\left(\max_{\theta\in P_{K}^{i}}{\langle\theta-\theta_{*}^{i},z_{k^{\prime}}\rangle}-z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})\right)$
    $+z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})$
    $\geq(-\alpha_{0})\langle\vartheta,z_{k^{\prime}}\rangle+(1+\alpha_{0})z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})$
    $\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta.$
  • If $\vartheta^{T}z_{k}\leq-\beta\delta$, by (10), we have
    $\max_{\theta\in P_{K}^{i}}{\langle\theta-g(P_{K}^{i}),z_{k^{\prime}}\rangle}\geq\frac{1}{n_{z}}\max_{\theta\in P_{K}^{i}}{\langle\theta-g(P_{K}^{i}),-z_{k^{\prime}}\rangle}.$ (57)
    Then, $\psi_{k^{\prime}}^{+}$ can be lower bounded by
    $\psi_{k^{\prime}}^{+}$ $\geq\frac{(-\alpha_{0})}{n_{z}}\max_{\theta\in P_{K}^{i}}{\langle\theta-g(P_{K}^{i}),-z_{k^{\prime}}\rangle}+z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})$
    $=\frac{(-\alpha_{0})}{n_{z}}\left(\max_{\theta\in P_{K}^{i}}{\langle\theta-\theta_{*}^{i},-z_{k^{\prime}}\rangle}+z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})\right)$
    $+z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})$
    $\geq\frac{(-\alpha_{0})}{n_{z}}\langle\vartheta,-z_{k^{\prime}}\rangle+\left(1+\frac{(-\alpha_{0})}{n_{z}}\right)z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})$
    $\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta.$
In both sub-cases, the desired conclusion holds. Therefore, regardless of the sign of $\vartheta^{T}z_{k^{\prime}}$, $\psi_{k^{\prime}}^{+}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta$ follows.
When $z_{k^{\prime}}^{T}(g(P_{K}^{i})-\theta_{*}^{i})<0$, a similar analysis can be applied to $\psi_{k^{\prime}}^{-}$ and we have
$\psi_{k^{\prime}}^{-}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta.$ (58)
Thus, in either case, we conclude that
$\max\{\psi_{k}^{+},\psi_{k}^{-}\}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta,$ (59)
which prove the claim (56).
Consider partitioning the time interval $[K+1:K+\Delta K]$ into a sequence of consecutive segments of length $N_{u}$. Define for $t=0,\dots,\Delta T-1$:
$\mathcal{I}_{t}\triangleq[K+tN_{u}+1:K+(t+1)N_{u}].$ (60)
If $\Delta K$ is divisible by $N_{u}$, define $\mathcal{I}_{\Delta T}\triangleq\emptyset$; otherwise, let
$\mathcal{I}_{\Delta T}\triangleq[K+\Delta TN_{u}+1:K+\Delta K].$ (61)
Then, we have
$[K+1:K+\Delta K]=\bigcup_{t=0}^{\Delta T-1}\mathcal{I}_{t}\cup\mathcal{I}_{\Delta T}.$ (62)
With this partition, by law of total probability, $\mathbb{P}(\mathcal{E}_{[K+1:K+\Delta K]})$ can be written as
$\mathbb{P}(\mathcal{E}_{[K+1:K+\Delta K]})=\prod_{t=0}^{\Delta T}\mathbb{P}(\mathcal{E}_{\mathcal{I}_{t}}\mid{\mathcal{E}_{[K+1:K+tN_{u}]}}),$ (63)
where, by convention, we write $\mathbb{P}(\mathcal{E}_{\mathcal{I}_{0}}\mid\mathcal{E}_{[K+1:K]})\triangleq\mathbb{P}(\mathcal{E}_{\mathcal{I}_{0}})$.
Fix $t\in[0:\Delta T-1]$, by Lemma 4.9, there exists $k_{t}\in\mathcal{I}_{t}$ such that $|\vartheta^{T}z_{k_{t}}|\geq\beta\delta$. We denote the following interval as $\mathcal{I}_{t,1}\triangleq[K+tN_{u}+1:k_{t}]$, $\mathcal{I}_{t,2}\triangleq[k_{t}+1:K+(t+1)N_{u}]$, $\mathcal{I}_{t,3}\triangleq[K+1:K+tN_{u}]$ and $\mathcal{I}_{t,4}\triangleq[K+1:k_{t}+1]$. Therefore, $\mathbb{P}(\mathcal{E}_{\mathcal{I}_{t}}\mid{\mathcal{E}_{[K+1:K+tN_{u}]}})$ can be further written as
$\mathbb{P}(\mathcal{E}_{\mathcal{I}_{t,1}}\mid{\mathcal{E}_{\mathcal{I}_{t,3}}})\mathbb{P}(\mathcal{E}_{k_{t}+1}\mid\mathcal{E}_{[K+1:k_{t}]})\mathbb{P}(\mathcal{E}_{\mathcal{I}_{t,2}}\mid\mathcal{E}_{\mathcal{I}_{t,4}}).$ (64)
Consider the second term in (64), $\mathcal{E}_{k_{t}+1}$ holds if and only if
$-\bar{w}(i)+\psi_{k_{t}}^{+}\leq w_{k_{t}}\leq\bar{w}(i)-\psi_{k_{t}}^{-}.$ (65)
With (56), $\mathbb{P}(\mathcal{E}_{k_{t}+1}\mid\mathcal{E}_{[K+1:k_{t}]})$ can be written as
$\mathbb{P}(\mathcal{E}_{k_{t}+1}\mid\psi_{k_{t}}^{+}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta,\mathcal{F}_{k_{t}})\mathbb{P}(\psi_{k_{t}}^{+}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta\mid\mathcal{F}_{k_{t}}))$
$+$ $\mathbb{P}(\mathcal{E}_{k_{t}+1}\mid\psi_{k_{t}}^{+}<\frac{(-\alpha_{0})\beta}{n_{z}}\delta,\mathcal{F}_{k_{t}})\mathbb{P}(\psi_{k_{t}}^{+}<\frac{(-\alpha_{0})\beta}{n_{z}}\delta\mid\mathcal{F}_{k_{t}}))$
$\leq$ $\left(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right)\right)\mathbb{P}(\psi_{k_{t}}^{+}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta\mid\mathcal{F}_{k_{t}}))$
$+$ $\left(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right)\right)\mathbb{P}(\psi_{k_{t}}^{+}<\frac{(-\alpha_{0})\beta}{n_{z}}\delta\mid\mathcal{F}_{k_{t}}))$
$=$ $\left(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right)\right).$
Therefore, we have
$\mathbb{P}(\mathcal{E}_{\mathcal{I}_{t}}\mid{\mathcal{E}_{[K+1:K+tN_{u}]}})\leq\left(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right)\right).$ (66)
Then, by (63), it follows that:
$\mathbb{P}(\mathcal{E}_{[K+1:K+\Delta K]})\leq\left(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right)\right)^{\Delta T}.$ (67)
Given any $p_{\varepsilon}>0$, if the right-hand side of (67) is less than $p_{\varepsilon}$, which holds when $\Delta T$ satisfies
$\Delta T>{\ln(p_{\varepsilon})}/{\ln(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right))}.$ (68)
Inequality (68) implies that the probability of $\mathcal{E}_{[K+1:K+\Delta K]}$ is no greater than $p_{\varepsilon}$ when $\Delta T$ exceeds the right-hand side of (68). Considering the complement event, with probability no less than $1-p_{\varepsilon}$, (46) holds.
For statement (iii), as $K$ tends to $\infty$, we show that $\forall i\in[n_{x}]$, $n^{i}(K)$ tends to $\infty$ almost surely. From previous analysis, the probability that $n^{i}(K)$ remains unchanged at each time step decays exponentially. This implies that for any fixed $i$, an increment of one occurs for $n^{i}(K)$ almost surely as $K$ increases. Therefore, as $K\to\infty$, $n^{i}(K)$ will almost surely increase without bound, which leads to $\lim_{K\to\infty}\mu(\widehat{\Theta}_{K})=0$.
Remark 4.15.
Theorem 4.13 discusses both non-asymptotic and asymptotic convergence of measure of feasible set under threshold trigger. For statement (i), an explicit upper bound of product measure of feasible set in terms of triggering count $n^{i}(K)$ is derived. For statement (ii), a non-asymptotic probability bound is formulated. Statement (ii) characterizes a probabilistic trigger interval for next uncertainty reduction. As scaler parameter $\alpha_{0}$ increases toward zero, the amount of uncertainty reduction per trigger increases in (45). But the triggering interval potentially becomes larger by (46), reflecting a trade-off between trigger frequency and convergence speed. Statement (iii) shows that the measure of uncertainty size is asymptotically convergent to 0 almost surely as $K\to\infty$.
The convergence of product measure of feasible set offers a clear and quantitative characterization of uncertainty reduction during the learning process. Another useful form of convergence is known as point-wise convergence, which characterizes the probability that each individual parameter point, other than the true parameter, is excluded from the feasible set. We observe that in the proof of Theorem 4.13, the sign of two expression $z_{k}^{T}\vartheta$ and $z_{k}^{T}(g(P_{k}^{i})-\theta_{*}^{i})$ are closely related to the classical arguments used in establishing point-wise convergence. By introducing an additional condition on the signal of two expressions, we can also derive a convergence result in the point-wise sense. The condition is formalized in the following assumption.
Assumption 5
Under the threshold trigger mechanism and Assumption 4, there exists $N_{G}>0$ such that for any $i\in[n_{x}]$ and $k_{0}>0$, there exists $k\in[k_{0}+1:k_{0}+N_{G}]$ satisfying:
$|z_{k}^{T}\vartheta|\geq\beta\delta\quad\text{and}\quad\text{sgn}(z_{k}^{T}\vartheta)=-\text{sgn}(z_{k}^{T}(g(P_{k}^{i})-\theta_{*}^{i})).$ (69)
Corollary 4.16.
Under assumptions of Theorem 4.13 and Assumption 5, it follows that
$\lim_{K\to\infty}\widehat{\Theta}_{K}=\Theta_{*}.$ (70)
Proof 4.17.
To prove (70), it suffices to prove the following statement: For any $i\in[n_{x}]$,
$\lim_{K\to\infty}\mathbb{P}(\theta\in P_{K}^{i})=0$ (71)
for all $\theta\in\mathbb{R}^{n_{z}}$ with $\left\lVert\theta-\theta_{*}^{i}\right\rVert_{2}=\delta>0$.
Denote $\vartheta=\theta-\theta_{*}^{i}$. Let event $\mathcal{E}_{k}^{\prime}$ be defined as
$\mathcal{E}_{k}^{\prime}\triangleq\{w_{k}(i):-\bar{w}(i)-z_{k}^{T}\vartheta\leq w_{k}(i)\leq\bar{w}(i)-z_{k}^{T}\vartheta\}.$ (72)
Moreover, for $S\subseteq\mathbb{Z}_{>0}$, we define
$\mathfrak{T}_{S}\triangleq\{k\in S\mid\gamma_{k,i}=1\}$ (73)
as the index set when sub-trigger $\gamma_{k,i}=1$ in set $S$. Let event $\Omega_{S}$ be defined as
$\Omega_{S}\triangleq\{\cap_{k\in\mathfrak{T}_{S}}\mathcal{E}_{k}^{\prime}\}.$ (74)
Then, the probability of $\theta\in P_{K}^{i}$ can be written as
$\mathbb{P}(\theta\in P_{K}^{i})=\mathbb{P}(\Omega_{[K]})$
$=$ $\left(\prod_{t=0}^{\Delta T-1}\mathbb{P}(\Omega_{[tN_{G}+1:(t+1)N_{G}]}\mid\Omega_{[tN_{G}]})\right)\mathbb{P}(\Omega_{[(\Delta T)N_{G}:K]}),$ (75)
where, by convention, we write $\mathbb{P}(\Omega_{[N_{G}]}\mid\Omega_{0})$ as $\mathbb{P}(\Omega_{[N_{G}]})$ and $\Delta T=\lfloor\frac{K}{N_{G}}\rfloor$.
We assume $k$ to be the index satisfying (69), i.e., $u_{k}$ is designed such that $|z_{k}^{T}\vartheta|\geq\beta\delta\quad\text{and}\quad\text{sgn}(z_{k}^{T}\vartheta)=-\text{sgn}(z_{k}^{T}(g(P_{k}^{i})-\theta_{*}^{i}))$. Then, $\Omega_{k}$ implies
$\gamma_{k,i}=0\quad\text{or}\quad(\gamma_{k,i}=1\text{ and }\mathcal{E}_{k}^{\prime}).$ (76)
We consider the case $z_{k}^{T}(g(P_{k}^{i})-\theta_{*}^{i})\leq 0$ (The case for $z_{k}^{T}(g(P_{k}^{i})-\theta_{*}^{i})\geq 0$ can be discussed similarly). By the choice of $k$, $z_{k}^{T}\vartheta\geq\beta\delta>0$. From (58), we have $\psi_{k}^{-}\geq\frac{(-\alpha_{0})\beta}{n_{z}}\delta$. We consider the two sub-cases based on value of $\psi_{k}^{-}$ and $z_{k}^{T}\vartheta$ and show that the following implication holds:
$\theta\in P_{k+1}^{i}\implies w_{k}(i)\leq\bar{w}_{i}-\frac{(-\alpha_{0})\beta}{n_{z}}\delta.$ (77)
  • $\psi_{k}^{-}\leq z_{k}^{T}\vartheta$: From (29), $\gamma_{k,i}=1$ implies $w_{k}(i)\geq\bar{w}_{i}-\psi_{k}^{-}$ or $w_{k}(i)\leq-\bar{w}_{i}+\psi_{k}^{+}$. If $\psi_{k}^{i}\leq z_{k}^{T}\vartheta$, $w_{k}(i)\geq\bar{w}_{i}-\psi_{k}^{-}$ contradicts with $\mathcal{E}_{k}^{\prime}$. Combining (76), we have $\theta\in P_{k+1}^{i}\implies\gamma_{k,i}=0$, which can be equivalently formulated as $-\bar{w}_{i}+\psi_{k}^{+}\leq w_{k}(i)\leq\bar{w}_{i}-\psi_{k}^{-}$. Recall that $\psi_{k}^{-}$ is lower bounded by (58), we have (77) holds.
  • $\psi_{k}^{-}>z_{k}^{T}\vartheta$: In this case, $\gamma_{k,i}=1$ and $\mathcal{E}_{k}^{\prime}$ together imply $\bar{w}(i)-\psi_{k}^{-}\leq w_{k}(i)\leq\bar{w}(i)-z_{k}^{T}\vartheta$ or $w_{k}(i)\leq-\bar{w}(i)+\psi_{k}^{+}$. The event $\gamma_{k,i}=0$ implies $-\bar{w}(i)+\psi_{k}^{+}\leq w_{k}(i)\leq\bar{w}(i)-\psi_{k}^{-}$. Thus, $w_{k}(i)$ satisfies $w_{k}(i)\leq\bar{w}(i)-z_{k}^{T}\vartheta$. Recall that $z_{k}^{T}\vartheta>\beta\delta>\frac{(-\alpha_{0})\beta}{n_{z}}\delta$, we conclude that (77) holds.
The above discussion shows that for any index $k$ satisfying the condition in (69), the event $\theta\in P_{k+1}^{i}$ implies that $w_{k}(i)$ satisfy $w_{k}(i)\leq\bar{w}_{i}-\frac{(-\alpha_{0})\beta}{n_{z}}\delta$ or $w_{k}(i)\geq-\bar{w}_{i}+\frac{(-\alpha_{0})\beta}{n_{z}}\delta$. Since $w_{k}(i)$ is independent of $z_{k}$ and the information $\mathcal{F}_{k}$ available at step $k$, the conditional probability of this event given $\mathcal{F}_{k}$ has the following expression
$\mathbb{P}(\Omega_{k}\mid\mathcal{F}_{k})\leq 1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right).$ (78)
Now we derive upper-bound for each $\mathbb{P}(\Omega_{[tN_{G}+1:(t+1)N_{G}]}\mid\Omega_{[tN_{G}]})$. Under Assumption 5, there exist $k\in[tN_{G}+1:(t+1)N_{G}]$ such that (69) holds. Then, we have
$\mathbb{P}(\Omega_{[tN_{G}+1:(t+1)N_{G}]}\mid\Omega_{[tN_{G}]})$
$=$ $\mathbb{P}(\Omega_{[tN_{G}:k-1]}\mid\Omega_{[tN_{G}]})\mathbb{P}(\Omega_{k}\mid\Omega_{[k-1]})$
$\mathbb{P}(\Omega_{[k+1:(t+1)N_{G}]}\mid\Omega_{[k]})$
$\leq$ $\mathbb{P}(\Omega_{k}\mid\Omega_{[k-1]})\leq 1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right).$
Therefore, the probability of $\theta\in P_{k}^{i}$ derived in (75) can be upper-bounded by
$\mathbb{P}(\theta\in P_{K}^{i})\leq\left(1-q\left(\frac{(-\alpha_{0})\beta}{n_{z}}\delta\right)\right)^{\lfloor\frac{K}{N_{G}}\rfloor}.$ (79)
From (79), (71) holds as the probability exponentially decreases to 0 as $K\to\infty$, which proves the Corollary 4.16.
Remark 4.18.
Corollary 4.16 provides a point-wise convergence for feasible set. This extended result is enabled by Assumption 5, which plays a key role analogous to persistent excitation. It ensures that, infinitely often, the system input $u_{k}$ together with $w_{k}$ generates sufficiently informative signal that the signs of two expressions does not be identical all the time when $z_{k}$ satisfies $|z_{k}^{T}\vartheta|>\beta\delta$. The structure of Assumption 5 also gives insights into input design of $u_{k}$. It highlights how appropriately chosen inputs $u_{k}$ can theoretically ensure not only reduction of uncertainty in measure, but also convergence of estimator to the true parameters.
4.2 Convergence under Immediate trigger $(\alpha_{0}=-1)$
In this section, we study the convergence behavior of the measure of the feasible set $\widehat{\Theta}_{K}$ under the immediate trigger condition, where the predefined threshold is set as $\alpha_{0}=-1$.
The following proposition captures a distinctive property specific to the case when $\alpha_{0}=-1$:
Proposition 4.19.
Consider system (1) under Assumptions 1 and 2. Then, the stacked polyhedral set under the immediate trigger condition ($\alpha_{0}=-1$) coincides with that constructed using the full dataset, i.e.,
$\widehat{\Theta}_{K}=\widehat{\Theta}_{K}^{\text{full}}.$ (80)
Proof 4.20.
To prove the equality of two sets, it suffices to show mutual inclusion: $\widehat{\Theta}_{K}^{\text{full}}\subseteq\widehat{\Theta}_{K}$ and $\widehat{\Theta}_{K}\subseteq\widehat{\Theta}_{K}^{\text{full}}$.
We first show that $\widehat{\Theta}_{K}^{\text{full}}\subseteq\widehat{\Theta}_{K}$. For each local set $P(\mathcal{D}_{[K]}^{i})$ in the construction of $\widehat{\Theta}_{K}^{\text{full}}$, consider all hyperplane constraints that define $P(\mathcal{D}_{[K]}^{i})$. Since the triggered dataset is a subset of the full dataset, all hyperplane constraints used in constructing $P_{K}^{i}$ also appear among those defining $P(\mathcal{D}_{[K]}^{i})$. Therefore, it holds that $P_{K}^{i}\subseteq P(\mathcal{D}_{[K]}^{i})$ for each $i$, and thus $\widehat{\Theta}_{K}^{\text{full}}\subseteq\widehat{\Theta}_{K}$.
Next, we prove $\widehat{\Theta}_{K}\subseteq\widehat{\Theta}_{K}^{\text{full}}$. We show that every hyperplane constraint appearing in the minimal representation of $\widehat{\Theta}_{K}$ is also included in the representation of $\widehat{\Theta}_{K}^{\text{full}}$. This follows because, when a constraint generated from $\mathcal{D}_{k}$ is necessary for the minimal description, there exists an $i\in[n_{x}]$ such that $P(\mathcal{D}_{k}^{i})$ intersects nontrivially with $P_{k-1}^{i}$. Consequently, $\gamma_{k,i}=1$ by the immediate trigger condition, and thus the constraint from $\mathcal{D}_{k}$ is incorporated into the construction of $\widehat{\Theta}_{S_{K}}$, ensuring $\widehat{\Theta}_{K}\subseteq\widehat{\Theta}_{K}^{\text{full}}$.
Therefore, we conclude $\widehat{\Theta}_{K}^{\text{full}}=\widehat{\Theta}_{K}$.
Remark 4.21.
Proposition 4.19 implies that, under the immediate trigger condition, all process data contributing to the boundary of polyhedral feasible set are guaranteed to be retained. As a result, the feasible set coincides with that constructed from the full dataset. This equivalence enables the direct application of existing convergence analysis for SME using full dataset to the case where the immediate trigger condition is in effect.
We now introduce a lemma that characterizes the size of each polyhedral set.
Lemma 4.22.
Let $q(\epsilon)$ be a lower bound function of visiting boundary property for $w$. For any $\delta>0,i\in[n_{x}]$, the probability of $\mathfrak{R}(P(\mathcal{D}_{K}^{i}),\theta_{*}^{i})>\delta$ is upper bounded by
$\mathbb{P}(\mathfrak{R}(P_{K}^{i}),\theta_{*}^{i})>\delta)\leq\bar{\mathcal{N}}\left(1-q(\frac{\beta\delta}{2})\right)^{\lfloor\frac{K}{N_{u}}\rfloor},$ (81)
where $\bar{\mathcal{N}}=(\frac{4b_{z}}{\beta}+1)^{n_{z}}$ is an upper bound of covering numbers $\mathcal{N}(B_{2}^{n_{z}},\frac{\beta}{2b_{z}})$ of the Euclidean ball $B_{2}^{n_{z}}$.
Proof 4.23.
A detailed proof is omitted here for brevity and can be reconstructed following similar steps in [31].
Remark 4.24.
Lemma 4.22 provides a probabilistic guarantee on the radius of each polyhedral set $P_{K}^{i}$. Specifically, it shows that $\forall i\in[n_{x}]$ and for a fixed $\delta>0$, as the number of samples $K$ increases, the probability of $\mathfrak{R}(P(\mathcal{D}_{K}^{i}),\theta_{*}^{i})>\delta$ decays to 00 exponentially. This result ensures eliminating uncertainty in parameter estimation.
Now we are in a position to derive the theorem for measure convergence of feasible set under immediate trigger selection condition.
Theorem 4.25.
Consider system (1) under the immediate trigger selection strategy $(\alpha_{0}=-1)$. Suppose that Assumptions 1–4 hold, and $q(\epsilon)$ be a lower bound function in (35). Then, the estimator $\widehat{\Theta}_{K}$ satisfies the following two properties.
  • (i)
    With probability no less than
    $1-n_{x}\bar{\mathcal{N}}\left(1-q\left(\frac{\beta\delta}{2}\right)\right)^{\lfloor K/N_{u}\rfloor},$ (82)
    it holds that
    $\mu(\widehat{\Theta}_{K})\leq\left(\delta^{n_{z}}\frac{\pi^{n_{z}/2}}{\Gamma\left(\frac{n_{z}}{2}+1\right)}\right)^{n_{x}}.$ (83)
  • (ii)
    Moreover, as $K\to\infty$, the following convergence holds almost surely:
    $\lim_{K\to\infty}\mu(\widehat{\Theta}_{K})=0.$ (84)
Proof 4.26.
The proof is divided into two parts, corresponding to statements (i) and (ii).
For the statement (i), we first prove that the following probability bound holds:
$\mathbb{P}(\max_{i\in[n_{x}]}\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})>\delta))\leq{n_{x}}\bar{\mathcal{N}}\left(1-q(\frac{\beta\delta}{2})\right)^{\lfloor\frac{K}{N_{u}}\rfloor}$ (85)
The event $\max_{i\in[n_{x}]}\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})>\delta$ imples that there exists $i\in[n_{x}]$ such that $\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})>\delta$. Consequently, by the union bound, we have the following upper bound:
$\mathbb{P}(\max_{i\in[n_{x}]}\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})>\delta))\leq\sum_{i=1}^{n_{x}}\mathbb{P}(\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})>\delta).$ (86)
Here we have used the fact that the event $\{\max_{i\in[n_{x}]}\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})>\delta\}$ is contained within the union of events $\cup\{\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})>\delta\}$. Then, by Lemma 4.22, the inequality (85) holds.
If we have $\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})\leq\delta$, it means that every point $\theta\in P_{K}^{i}$ lies within a closed ball $\mathcal{B}_{n_{z}}(\theta_{*}^{i},\delta)$ centered at $\theta_{*}^{i}$ with radius $\delta$. This ball is a hypersphere with diameter $2\delta$. The volume of an $n_{z}$-dimensional hypersphere with radius $\delta$ is given by
$\mu(\mathcal{B}_{n_{z}}(\theta_{*}^{i},\delta))=\frac{\pi^{\frac{n_{z}}{2}}}{\Gamma(\frac{n_{z}}{2}+1)}\delta^{n_{z}},$ (87)
where $\Gamma(\cdot)$ is Euler’s gamma function. Then, $P_{K}^{i}\subseteq\mathcal{B}_{n_{z}}(\theta_{*}^{i},\delta)$ implies that
$\mu(P_{K}^{i})\leq\mu(\mathcal{B}_{n_{z}}(\theta_{*}^{i},\delta)).$ (88)
From the above discussion, we show that with probability at least $1-{n_{x}}\bar{\mathcal{N}}\left(1-q(\frac{\beta\delta}{2})\right)^{\lfloor\frac{K}{N_{u}}\rfloor}$, $\max_{i\in[n_{x}]}\mathfrak{R}(P_{K}^{i},\theta_{*}^{i})\leq\delta$, then all $P_{K}^{i}(i\in[n_{x}])$ are included in $n_{z}$-dimensional hypersphere whose radius are $\delta$. Then, by definition of product measure of $\widehat{\Theta}_{K}$ in (41), we have
$\mu(\widehat{\Theta}_{K})\leq\prod_{i=1}^{n_{x}}\mu(\mathcal{B}_{n_{z}}(\theta_{*}^{i},\delta))=\left(\delta^{n_{z}}\frac{\pi^{\frac{n_{z}}{2}}}{\Gamma(\frac{n_{z}}{2}+1)}\right)^{n_{x}}$ (89)
which completes the proof of statement (i).
For Statement (ii), we aim to show that $\mu(\widehat{\Theta}_{K})$ converges to zero in probability as $K\to\infty$.
Fix a $\delta>0$. From (83), the upper bound of $\mu(\widehat{\Theta}_{K})$ can be expressed as $c\cdot\delta^{n_{z}n_{x}}$, where $c>0$ is a constant independent of $\delta$. Choose $\delta>0$ such that $c\cdot\delta^{n_{z}n_{x}}=\varepsilon$. By Lemma 4.6, we have
$q\left(\frac{\beta\delta}{2}\right)>0\quad\text{for all}\quad\delta>0.$
Therefore, it follows that
$\lim_{K\to\infty}n_{x}\bar{\mathcal{N}}\left(1-q\left(\frac{\beta\delta}{2}\right)\right)^{\lfloor K/N_{u}\rfloor}=0.$ (90)
Hence, for any $\varepsilon>0$, the probability that $\mu(\widehat{\Theta}_{K})\leq\varepsilon$ tends to 1 as $K\to\infty$, which establishes Statement (ii).
Remark 4.27.
Theorem 4.25 establishes both non-asymptotic and asymptotic convergence results for the measure of the estimator $\widehat{\Theta}_{K}$ under the immediate trigger selection strategy. In particular, statement (i) provides a finite-sample probabilistic bound for the measure of the feasible set, while statement (ii) guarantees the almost sure convergence of $\mu(\widehat{\Theta}_{K})$ to zero as $K$ tends to infinity. Moreover, from (82), the convergence rate depends on the properties of the function $q(\cdot)$. For a fixed radius $\delta$, a larger value of $q(\delta)$ leads to a faster convergence rate, which corresponds to the case where the random disturbance exhibits a higher probability measure near the boundary of its support. For instance, given a compact and convex set $\mathcal{W}$, the convergence of the estimator under uniformly distributed $w_{k}$ is faster than that under Gaussian distributed $w_{k}$.
4.3 Selection Probability Bound
This section investigates the statistical properties of the coreset selection mechanism, focusing on the selection probability and the cumulative number of selected data points. By triggering condition (30) and Proposition 4.19, the selection probability associated with any $\alpha_{0}\in(-1,0)$ is always upper bounded by the case of immediate trigger. We derive one theoretical property on the selection probability upper bound under immediate trigger. The following theorem establishes such a bound.
Theorem 4.28.
Let Assumptions 14 be satisfied. $q$ and $Q$ are coordinate-wise boundary function for $w_{k}$. Under immediate trigger, for any $\delta>0$, the expectation of selection probability at time $k$ is upper bounded by the following inequalities:
$\mathbb{E}[\mathbb{P}(\gamma_{k}=1\mid\mathcal{F}_{k-1})]\leq n_{x}[\bar{\mathcal{N}}\left(1-q(\frac{\beta\delta}{2})\right)^{\lfloor\frac{k}{N_{u}}\rfloor}+Q(b_{z}\delta)].$ (91)
Proof 4.29.
To establish the validity of (91), we introduce reference ball $B_{n_{z}}(\theta_{*}^{i},{\delta})$ centered at $\theta_{*}^{i}$ and evaluating the intersection probability of new hyperplane constraints with $\widehat{\Theta}_{k}$ through probabilistic analysis.
The expectation of selection probability can be written as
$\mathbb{E}[\mathbb{P}(\gamma_{k}=1\mid\mathcal{F}_{k-1})]=\mathbb{P}_{w_{1},...,w_{k}}(\gamma_{k}=1).$
We denote the $i$-th new feasible set as
$\Delta_{k}^{i}=H_{+}^{i}\cap H_{-}^{i},$ (92)
where $H_{+}^{i}$ and $H_{-}^{i}$ are defined in (23) and (24). Let $\mathcal{J}_{k}^{i}$ denotes the event $\{P_{k-1}^{i}\nsubseteq\Delta_{k}^{i}\}$. Under immediate trigger, we have $\gamma_{k,i}=1\iff\mathcal{J}_{k}^{i}$. Then, $\gamma_{k}=1$ implies that $\cup_{i=1}^{n_{x}}\mathcal{J}_{k}^{i}$ is true. Thus, the following inequality holds:
$\mathbb{P}_{w_{1},...,w_{k}}(\gamma_{k}=1)\leq\sum_{i=1}^{n_{x}}\mathbb{P}_{w_{1},...,w_{k}}(\mathcal{J}_{k}^{i}).$ (93)
Let $\mathcal{B}^{i}_{\delta}$ denote the event $\{\mathfrak{R}(P_{k-1}^{i},\theta_{*}^{i}))\geq\delta_{k}\}$ and $\overline{\mathcal{B}^{i}_{\delta}}$ be its complement $\{\mathfrak{R}(P_{k-1}^{i},\theta_{*}^{i}))<\delta_{k}\}$. For a fixed $i\in[n_{x}]$, By law of total probability, we have that
$\mathbb{P}(\mathcal{J}_{k}^{i})=\mathbb{P}(\mathcal{J}_{k}^{i}\mid\mathcal{B}^{i}_{\delta})\mathbb{P}(\mathcal{B}^{i}_{\delta})+\mathbb{P}(\mathcal{J}_{k}^{i}\mid\bar{\mathcal{B}^{i}_{\delta}})\mathbb{P}(\bar{\mathcal{B}^{i}_{\delta}})$ (94)
where $\mathbb{P}$ is on full probability space $\{w_{1},...,w_{k}\}$. Since the probability measure $\mathbb{P}(\cdot)$ is strictly bounded above by 1, (94) can be simplified to
$\mathbb{P}(\mathcal{J}_{k}^{i})\leq\mathbb{P}(\mathcal{B}^{i}_{\delta})+\mathbb{P}(\mathcal{J}_{k}^{i}\mid\bar{\mathcal{B}^{i}_{\delta}}).$ (95)
The first term can be upper bounded by Lemma 4.22. To get an upper bound for the second term, we define
$h_{+}^{i}$$\triangleq{\{\theta:x_{k+1}(i)-z_{k}^{T}\theta=\bar{w}(i)\}},$ (96)
$h_{-}^{i}$$\triangleq{\{\theta:x_{k+1}(i)-z_{k}^{T}\theta=-\bar{w}(i)\}}.$ (97)
Let $d(h,\theta_{*}^{i})$ denote the Euclidean distance from the point $\theta_{*}^{i}$ to the hyperplane $h$. Also, we define $d(h^{i},\theta_{*}^{i})=\min\{d(h_{+}^{i},\theta_{*}^{i}),d(h_{-}^{i},\theta_{*}^{i})\}$.
First, we prove that the probability of $d(h^{i},\theta_{*}^{i})<\delta$ is upper bounded by $Q(b_{z}\delta)$. From the definition of $d$, we have
$d(h_{+}^{i},\theta_{*}^{i})=\frac{|w_{k}(i)-\bar{w}(i)|}{\|z_{k}\|_{2}},$ (98)
and
$d(h_{-}^{i},\theta_{*}^{i})=\frac{|w_{k}(i)+\bar{w}(i)|}{\|z_{k}\|_{2}}.$ (99)
Thus, $d(h^{i},\theta_{*}^{i})<\delta$ implies $\bar{w}(i)-|w_{k}(i)|<\delta\|z_{k}\|_{2}\leq b_{z}\delta$. From Lemma 4.6: $\mathbb{P}(d(h^{i},\theta_{*}^{i})<\delta)\leq Q(b_{z}\delta)$ holds for all $k>0$ and $i\in[n_{x}]$.
Given $\overline{\mathcal{B}^{i}_{\delta}}$ holds, if event $\mathcal{J}_{k}^{i}$ is true, the minimum distance between $\theta_{*}^{i}$ and new hyperplane $h_{+}^{i},h_{-}^{i}$ is less than $\delta$. We have
$\mathbb{P}_{w_{1},...,w_{n}}(\mathbb{P}(\mathcal{J}_{k}^{i}\mid\overline{\mathcal{B}^{i}_{\delta}}))$ $\leq\mathbb{P}_{w_{1},...,w_{n}}(d(h^{i},\theta_{*}^{i})<\delta)$ (100)
$=\mathbb{P}_{w_{k}}(d(h^{i},\theta_{*}^{i})<\delta)$ (101)
$\leq Q(b_{z}\delta).$ (102)
By substituting the established upper bounds for the first and second terms in equation (94), and integrating these results with the inequality in equation (93), we conclude inequality (91).
Remark 4.30.
Theorem 4.28 provides an upper bound for expected selection probability for all $\alpha_{0}\in[-1,0)$. It’s worth mention that this bound holds for all $\delta$ in statement of Theorem 4.28. The optimal probability upper bound can be derived by minimizing the right-hand side of (91) with respect to $\delta$.
Let the cumulative number of selected data points over time horizon $[K]$ be denoted by
$\mathcal{T}(K)\triangleq\sum_{k=1}^{K}\mathbf{1}\{\gamma_{k}=1\}.$ (103)
If $q$ and $Q$ satisfy the following assumption:
Assumption 6
The boundary functions $q(\epsilon)$ and $Q(\epsilon)$ admit the following local expansion:
$q(\epsilon)=C_{1}\epsilon^{p}+O(\epsilon^{p+p^{\prime}}),Q(\epsilon)=C_{2}\epsilon^{p}+O(\epsilon^{p+p^{\prime}})$ (104)
for some $C_{1}$,$C_{2}>0$,$p>0$, and $p^{\prime}>0$.
Then, we have the following upper bound on the expected cumulative number of selected data points, which implies a vanishing average selecting rate.
Corollary 4.31.
Let Assumptions 14 and Assumption 6 hold. The expected cumulative number of selected data points satisfies
$\mathbb{E}(\mathcal{T}(K))\leq O((\ln K)^{2}).$ (105)
Proof 4.32.
The proof is provided in Appendix 8.
Remark 4.33.
Assumption 6 is satisfied by a broad class of disturbance distributions, including uniform distributions and truncated Gaussian distributions over hypercubes (cite for [31]). Corollary 4.31 establish a convergent bound for cumulative number of selected data. While the exact constants in the $O((\ln K)^{2})$ bound are not specified, the result captures the asymptotic scaling behavior w.r.t. $K$.
5 Simulations
In this section we provide simulation results to validate the theoretical results and show the effectiveness of the proposed online coreset selection strategy. We first present an illustrative example on a second-order system to visualize the uncertainty evolution and selection behavior.
5.1 Illustrative Example: Second-Order System
We begin by illustrating the proposed data selection method on a second-order discrete-time linear system of the form (1) with system matrices
$A=\begin{bmatrix}0.5366&0.2038\\ -0.0406&0.6310\end{bmatrix},\quad B=\begin{bmatrix}-0.02\\ 0.4730\end{bmatrix},$ (106)
as used in [36]. The disturbance sequence $w_{k}$ is i.i.d. and uniformly distributed over the ellipsoid set $\mathcal{W}=\{w\in\mathbb{R}^{2}:\|w\|_{2}\leq 0.5\}$. The trigger threshold parameter is set as $\alpha_{0}=-0.3$. And the total simulation length is $K=150$ steps. We assume that close-loop controller $u_{k}$ for sending agent has been appropriated design such that Assumption 4 holds, with $N_{u}=20$ and $\beta^{2}=0.05$. We choose the initial feasible set of $\theta_{0}^{i}(i\in\{1,2\})$ as a unit cube $\{x\in\mathbb{R}^{3}:\|x\|_{\infty}\leq 1\}$. Then it has $\mathcal{H}$-representation $P_{\mathcal{H}}([I;-I],\mathbf{1}_{6})$ and $\mathcal{V}$-representation $P_{\mathcal{V}}(V_{0})$ where columns of $V_{0}$ are vertices of unit cube.
5.1.1 Verification of PE
Fig. 2 represents the smallest eigenvalue of expression $\lambda_{\min}=(\sum z_{k}z_{k}^{T})/N_{u}$ over sliding windows. Specifically, each blue marker represents the minimum eigenvalue of the matrix $(\sum_{k=k_{0}-N_{u}+1}^{k_{0}}z_{k}z_{k}^{T})/N_{u}$ at time step $k_{0}$, while red dashed line denotes the constant threshold $\beta^{2}$. As observed, all eigenvalues remain above the threshold, thereby showing that the Assumption 4 is satisfied.
Figure 2: Verification of PE condition in Assumption 4: minimum eigenvalue of the covariance matrix over sliding windows of length $N_{u}=20$.
5.1.2 Data Selection and Uncertainty Reduction
To illustrate the effect of threshold-triggered data distillation, we simulate the evolution of the feasible set over time. The result is shown in Fig. 3 and Fig. 4.
Fig. 3 shows the evolution of measure $\mu(\widehat{\Theta}_{k})$ and cumulative number of selected data points $\mathcal{T}(k)$. From Fig. 3(a), we see that the measure of feasible set decreases steadily as more informative data is selected. From Fig. 3(b), we see that data selection is triggered only 29 times over the entire 150-step horizon. Together, Figs. 3(a) and (b) illustrate a representative execution of the proposed online coreset selection method.
Figure 3: Performance of the threshold-triggered estimator ($\alpha_{0}=-0.3$). (a) Evolution of the measure $\mu(\widehat{\Theta}_{k})$. (b) Cumulative number of selected data points.
Fig. 4 visualizes the uncertainty contraction process. The polyhedral set $P_{k}^{1}$ and $P_{k}^{2}$ are shown at time steps $k=0$, $k=20$, and $k=80$. The true parameter values are marked by red stars. It is observed that the volume of each 3D polytope reduces significantly over time, and the sets progressively contract toward the true parameter values, validating the effectiveness of the proposed data selection strategy in selecting informative data.
Figure 4: Evolution of feasible parameter sets. The top and bottom rows depict the 3D polyhedral uncertainty sets $P^{1}_{k}$ and $P^{2}_{k}$ at $k=0$, $k=20$, and $k=80$, respectively. Red stars mark the true parameter values.
5.2 Performance Under Different Disturbances
In this section we simulate the performance of the proposed coreset selection scheme under different disturbance distributions. Specifically, we compare two cases: truncated Gaussian disturbance and uniformly distributed disturbance.
We consider the linearized discrete-time system of Boeing 747 [31] of the form (1) with the following system matrices:
$A=\begin{bmatrix}0.99&0.03&-0.02&-0.32\\ 0.01&0.47&4.7&0\\ 0.02&-0.06&0.4&0\\ 0.01&-0.04&0.72&0.99\end{bmatrix},B=\begin{bmatrix}0.01&0.99\\ -3.44&1.66\\ -0.83&0.44\\ -0.47&0.25\end{bmatrix}.$ (107)
In the first case the disturbance sequence $\{w_{k}\}$ are sampled from Truncated Gaussian distribution generated by Gaussian distribution with zero mean and 1 variance with truncated range $[-2,2]$. In the second case, the disturbance is sampled from uniform distributed hypercube $\mathcal{W}=\{w\in\mathbb{R}^{4}:\|w\|_{\infty}\leq 2\}$.
The threshold parameter is set $\alpha_{0}=-0.3$, and the simulation horizon is set $K=150$. The controller is assumed to generate a regressor sequence satisfying Assumption 4 with $N_{u}=20$ and $\beta^{2}=0.02$. The initial feasible set of $\theta_{0}^{i}(i\in\{1,2,3,4\})$ are chosen as a hypercube $\{x\in\mathbb{R}^{6}:\|x\|_{\infty}\leq 10\}$, with $\mathcal{H}$-representation $P_{\mathcal{H}}([I;-I],10\times\mathbf{1}_{12})$ and $\mathcal{V}$-representation $P_{\mathcal{V}}(V_{0})$, where columns of $V_{0}$ are vertices of the hypercube with edge length 20 centered at the origin.
We conduct $N=100$ Monte Carlo trials. For each trail, regressor is repeatedly sampled until it satisfies the PE condition. The evolution of feasible set measure and the cumulative number of selected data points are recorded at each epoch.
Fig. 5 represents the results of all trials. We observe from Fig. 5(a) that truncated Gaussian disturbance leads to a consistently larger feasible set measure compared to the uniform case. From Fig. 5(b) we observe that the selection count is generally lower in truncated case. Fig. 5(a) and Fig. 5(b) together show effectiveness of the proposed selection trigger design for both disturbance scenarios.
Figure 5: Comparison of performance under truncated Gaussian and uniform disturbances. (a) Evolution of feasible set measure $\mu(\widehat{\Theta}_{k})$. (b) Cumulative number of selected data points. Solid lines indicate the average across 100 trials; The shaded regions represent the range between the minimum and maximum values.
5.3 Performance Under Different Scaling Parameters $\alpha_{0}$
We now examine how the parameter $\alpha_{0}$ influence both the reduction in parameter uncertainty and size of coreset.
The system for simulation is the same as (107), where we adopt uniform distributed disturbance. We vary the trigger parameter over $\alpha_{0}\in\{-1,-0.5,-0.3,-0.2,0\}$.
We randomly generate $N_{epoch}=100$ valid regressor sequences ${z_{k}}$ satisfying Assumption 4. For each trial, we record the evolution of the feasible set measure and the cumulative number of selected data points. The averaged results are shown in Fig. 6.
Figure 6: Comparison of performance under different scaling parameter $\alpha_{0}\in\{-1,-0.5,-0.3,-0.2,0\}$. (a) Evolution of feasible set measure $\mu(\widehat{\Theta}_{k})$. (b) Cumulative number of selected data points under different $\alpha_{0}$ values. The shaded regions represent the range between the minimum and maximum values.
As shown in Fig. 6(a), increasing threshold $\alpha_{0}$ enlarges the average feasible set, indicating higher uncertainty. Notably, the case $\alpha_{0}=-1$ corresponds to immediate trigger and yields a feasible set equivalent to that derived from the full dataset (cf. Proposition 4.19), serving as a baseline for comparison. Fig. 6(b) illustrates the cumulative number of selection with different $\alpha_{0}$. For instance, at $\alpha_{0}=-0.5$, the coreset size is around 53 over 500 steps, achieving a 90% reduction relative to full dataset, while maintaining uncertainty close to the baseline. This highlights the efficiency of the proposed threshold trigger mechanism for data selection.
6 Conclusion and Future Directions
In this work, we propose an online coreset selection method for set-membership identification with guaranteed convergence. In the proposed design, first an over-approximation of feasible set in the form of stacked polyhedral set is derived. Leveraging this representation, new data are kept when its corresponding hyperplane constraint has a scaling parameter w.r.t. current uncertainty larger than predefined threshold. The measure convergence for threshold trigger is proved with generalized Grunbaum’s inequality and persistently exciting data and tight disturbance bound. The case for immediate trigger is proved to be equivalent to cases with full dataset. We also derive an expected upper bound for number of data points in coreset. Finally, we conduct numerical simulations and validate the theoretic results. Future work will focus on extending the proposed framework to more general system classes and relaxing key assumptions, such as adopting probabilistic excitation conditions like the BMSB condition.
[…]
7 Proof of Lemma 4.6
Proof 7.34.
Let $\mathbb{P}_{i}$ denote the $i$-the marginal measure on $\pi_{i}(\mathcal{W})=[-\bar{w}(i),\bar{w}(i)]$. Let $M_{i,\epsilon}\triangleq\{w_{s}\in[-\bar{w}(i),\bar{w}(i)]:|w_{s}-\bar{w}(i)|<\epsilon\}$. For $\epsilon\leq\bar{w}(i)$, this is the interval $(\bar{w}(i)-\epsilon,\bar{w}(i)]$. Let $S_{i,\epsilon}=\{w\in\mathcal{W}:w(i)\in(\bar{w}(i)-\epsilon,\bar{w}(i)]\}$. Then $\mathbb{P}_{i}(M_{i,\epsilon})=\mathbb{P}(S_{i,\epsilon})$.
We define $q(\epsilon)=\inf_{i\in[n_{x}]}\mathbb{P}(S_{i,\epsilon})$. We show that $q$ satisfy the properties in Lemma 4.6. Clearly, $q(0)=0$. Since $S_{i,\epsilon}\subseteq S_{i,\epsilon^{\prime}}$ for $\epsilon\leq\epsilon^{\prime}$, $q$ is non-decreasing. Now we show that $q(\epsilon)>0$ holds for any $\epsilon>0$. For $i\in[n_{x}]$, choose $\tilde{w}\in\partial\mathcal{W}$ such that $\tilde{w}(i)=\bar{w}(i)$. The ball $B(\tilde{w},\epsilon)=\{w^{\prime}\in\mathcal{W}:||w^{\prime}-\tilde{w}||_{2}<\epsilon\}$ is contained in $S_{i,\epsilon}$, since any $w^{\prime}\in B(\tilde{w},\epsilon)$ satisfies $|w^{\prime}(i)-\bar{w}(i)|=|w^{\prime}(i)-\tilde{w}(i)|\leq||w^{\prime}-\tilde{w}||_{2}<\epsilon$, which implies $w^{\prime}(i)>\bar{w}(i)-\epsilon$. Since $w^{\prime}(i)\leq\bar{w}(i)$, $w^{\prime}(i)\in(\bar{w}(i)-\epsilon,\bar{w}(i)]$. By Assumption 3, $\mathbb{P}(B(\tilde{w},\epsilon))\geq q_{w}(\epsilon)$. Thus, $\mathbb{P}(S_{i,\epsilon})\geq\mathbb{P}(B(\tilde{w},\epsilon))\geq q_{w}(\epsilon)$. It follows that $q(\epsilon)=\inf_{i}\mathbb{P}(S_{i,\epsilon})\geq q_{w}(\epsilon)$. Since $q_{w}(\epsilon)>0$ for $\epsilon>0$, $q(\epsilon)>0$.
We define $Q(\epsilon)=\sup_{i\in[n_{x}]}\mathbb{P}(S_{i,\epsilon})$ for $\epsilon>0$. We show that $Q(0)=0$ and $Q$ are non-decreasing. Since $M_{i,0}$ is a single point or empty set, $\mathbb{P}(M_{i,0})=0$, so $Q(0)=0$. For $\epsilon_{1}leq\epsilon_{2}$, $M_{j,\epsilon_{1}}\subseteq M_{j,\epsilon_{2}}$. Thus, $\mathbb{P}(M_{j,\epsilon_{1}})\leq\mathbb{P}(M_{j,\epsilon_{2}})$. We have $Q(\epsilon_{1})<Q(\epsilon_{2})$. So $Q$ satisfies the property in Lemma 4.6.
8 Proof of Corollary 4.31
Proof 8.35.
Let $\varrho=\lfloor\frac{k}{N_{u}}\rfloor$ and $\lambda=\delta^{p}$. For small $\delta>0$, we have $q(\frac{\beta\delta}{2})=c_{1}\lambda+O(\lambda^{1+p^{\prime}/p}),Q(b_{z}\delta)=c_{2}\lambda+O(\lambda^{1+p^{\prime}/p})$. Then, (91) can be written as
$p(\lambda,\varrho)=c_{3}\left(1-c_{1}\lambda+O(\lambda^{1+p^{\prime}/p})\right)^{\varrho}+c_{4}\lambda+O(\lambda^{1+p^{\prime}/p}),$ (108)
where $c_{3}=n_{x}\bar{\mathcal{N}}$ and $c_{4}=n_{x}c_{2}$. We choose $\lambda(\varrho)=\frac{1}{c_{1}}(1-(\frac{c_{4}}{c_{1}\varrho})^{\frac{1}{\varrho-1}})$, which is the minimizer of function $p_{1}(\lambda)=c_{3}((1-c_{1}\lambda)^{\varrho}+c_{4}\lambda)$.
First we notice that the following relationship holds: $\lambda(\varrho)=\frac{\ln\varrho}{c_{1}\varrho}+o\left(\frac{\ln\varrho}{\varrho}\right)$. Moreover, we have that
$\lim_{\varrho\to\infty}\left(\frac{1-c_{1}\lambda+O(\lambda^{1+p^{\prime}/p})}{1-c_{1}\lambda}\right)^{\varrho}=1.$ (109)
Then, the probability can be reformulated as
$p(\lambda(\varrho),\varrho)=\frac{c_{4}\ln\varrho}{c_{1}\varrho}+o\left(\frac{\ln\varrho}{\varrho}\right).$ (110)
Therefore, the expected number of cumulative selected data points can be written in the following inequality.
$\mathbb{E}[\mathfrak{T}(K)]$ $=\sum_{k=1}^{K}\mathbb{E}[\mathbb{P}(\gamma_{k}=1)]$ (111)
$\leq O(\sum_{k=1}^{K}\frac{\ln k}{k})=O(\ln^{2}K),$ (112)
which completes the proof of Corollary 4.31.